A Complete Bibliography of The American Statistician: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.10

Title word cross-reference

$105.95 [Wel21]. $110.00 [CC24]. $120.00 [CF22, Mar24, Mic24, Ste22]. $122.95 [Bar22]. $130.00 [MV20]. $140.00 [Che21]. $159.99 [Per22a]. $180.00 [Ni24]. $250.00 [Par23]. $28.00 [Fle22, Knu21]. $29.95 [Wel20]. $32.00 [Boo21, Par20]. $32.95 [Kim20b, Per22b]. $38.95 [Joh20]. $54.95 [Ben21]. $59.95 [Kim23]. $64.99 [FA20]. $79.95 [Ben21, BS20b, Lun20]. $79.99 [Shi20]. $99.95 [Lee21, Rot22, Sos24, Wal24, You21]. D [GFB^{+}20]. e [Coh24]. F [Chr22b, GH22a, GH22b, Har22b]. M [GFY22]. P [ZSB22, And23, BB19, Fou20, Kli22, Pes20, SY21]. r [vdHZ22]. R^{2} [HWM24]. ρ [vdHZ22]. t [Gra23, GLW20]. τ [vdHZ22]. x + y [Knu21]. X/(X + Y) [Arn20]. Y/(X + Y) [Arn20].

-estimator [GFY22]. -Test [Gra23, Chr22b, GH22b, Har22b, GH22a]. -Tests [GLW20]. -Value [SY21, GFB^{+}20]. -Values [BB19, Fou20, Kli22, Pes20, ZSB22].

/2021 [Lee21].

0 [Che21, CC24, Mic24, Par20, Par23].

1 [Wel20]. 19 [CTÁM22, LBV^{+}22, Li22].

2 [Boo21, Fle22]. 2020 [JN24]. 29-36 [HJQ20b]. 2nd [Bar22, Che23, Cin23, Mic24, Sos24].

3 [Bar22]. 30-Day [PS21]. 3rd [CF22, Ni22].

4 [MV20, Per22a, Rot22]. 4th [Fle21b].

5 [FA20, Mar24, Per22b, Shi20].

6 [Kim20b, Knu21, Lun20, Wal24, Wel21].

7 [CF22, Kim23]. 73 [Fou20]. 74 [KX20, MH21, Tar20]. 75 [VW21].

8 [Joh20, Sos24, Ste22].

9 [BS20b, Ni24]. 978 [Bar22, Boo21, BS20b, Che21, CC24, CF22, Fle22, FA20, Joh20, Kim20b, Kim23, Knu21, Lun20, MV20, Mar24, Mic24, Ni24, Par20, Par23, Per22a, Per22b, Rot22, Shi20, Sos24, Ste22, Wal24, Wel20, Wel21]. 978-0-19-884439-6 [Kim20b]. 978-0-19-886416-5 [Per22b]. 978-0-367-00043-1 [Wel20]. 978-0-367-14067-0 [Par23]. 978-0-367-22109-6 [Lun20]. 978-0-367-34267-8 [Ste22]. 978-0-367-43667-4 [Rot22]. 978-0-367-63431-5 [Mar24]. 978-0-429-06666-5 [Lee21]. 978-0-465-09462-2 [Boo21]. 978-1-032-00745-8 [Sos24]. 978-1-032-06708-7 [Kim23]. 978-1-032-10571-0 [CC24]. 978-1-032-20325-6 [Wal24]. 978-1-032-23138-9 [Ni24]. 978-1-032-28826-0 [Mic24]. 978-1-108-49420-5 [Shi20]. 978-1-108-70529-5 [FA20]. 978-1-119-23857-7 [CF22]. 978-1-119-51265-3 [Bar22]. 978-1-138-07922-9 [BS20b]. 978-1-138-35933-8 [Joh20]. 978-1-138-62643-0 [Che21]. 978-1-138-62691-1 [Ben21]. 978-1-315-21266-1 [Ben21]. 978-1-315-36663-0 [You21]. 978-1-4398-8538-3 [Lee21]. 978-1-4987-4531-4 [MV20]. 978-1-5416-1851-0 [Par20]. 978-1-5416-4650-6 [Knu21]. 978-1-5416-9947-2 [Fle22]. 978-1-58488-939-7 [You21]. 978-1-63526-135-6 [Wel21]. 978-3-030-60059-4 [Per22a].

A/B [LSS^{+}24, SH22b]. Abramovich [Sos24]. Absence [PS21]. Accessible [HJB21]. Account [SH22b]. acov [Sos24]. Across [MKR^{+}23, HSHM^{+}23].

Activity [Gel23]. Actually [Hoa20]. Adaptations [HB22a]. Adjusting [LD21]. Adrian [Shi20]. Advanced [You21]. Adversarial [JNCP24, LCHL22]. Adversaries [GOIRS21]. Aekyung [KX20]. Agent [BH21, LBV^{+}22]. Agent-Based [BH21]. Airbnb [GH23]. Aitkin [CC24]. al [Chr22a, Lip23]. al. [RL23]. Algorithms [BYH^{+}23]. Aliasing [KM20]. Allaire [Joh20]. Allocation [KMG^{+}21]. Almost [CMB20]. Alpha [Bil22]. Alternating [Wal20]. Alternatives [McI22]. Amanda [Per22a]. American [Fou20, HJQ20b, KX20, MH21, Sch23, Tar20, VW21, Ano24, NYM24]. Analyses [HAT24, YR24]. Analysis [AG24, CTÁM22, wCW24, Che23, EIS20, HYW23, JNCP24, KM20, Li22, LAMM22, Lun20, Ni22, PS21, Ram23, Wal24, XC23]. Analytical [dMSNK23]. Analyzing [Kva21, NYM24]. Angel [Ni22]. Anomalies [ZZZ21]. ANOVA [Wes23]. Ante [SRSK22]. Application [Ano22, MFD^{+}23, MG24, Ngu21, YK21, YRD22]. Applications [HSLY23, KS21, SK20, Shi20, VH24, XGH^{+}20, Wel21]. Applied [Lee21, Che21, Mar24]. Approach [BP23, BS20b, HCS24, KBS22, LJ20, Lun20, Ngu21, Rot22, WSCP22, Wal24]. Approaches [Wan23]. Approximation [GdL23, O’N21]. Arbitrarily [Hua23]. Arbitrary [LLZL23]. Archetypal [EIS20]. Area [NCK^{+}22]. Army [JDC20]. Art [Par20]. Assessing [FG20]. Assessment [RUCH23]. Assignment [AB23]. Assignment-Control [AB23]. Associate [Ano24]. Associated [Ber21, MH20, MH21]. Association [YRD22]. Associations [vdHZ22]. Assumption [BBB^{+}21]. Assumptions [GFB^{+}20, SLS22]. Astros [EM22]. Asymmetric [HY23]. Athlete [MWS23]. Attack [LCHL22]. Attain [GHK24]. Attention [GBB20]. Authentic [Fre21]. Authors [HSHM^{+}23]. Autoregressive [Lar24]. Average [AM23, Gal20, GHK24]. Averaging [KBSM23].

B [LSS^{+}24, SH22b]. Bad [And23]. Bang [EM22]. Barcelona [PG23]. Bartroff [Sch23]. Baseball [GWJ22]. Based [BH21, BKL^{+}22, Chr22a, EIS20, GZWL24, HDDOV22, Hit24, Kli22, Lip23, WLL20, dMSNK23, CCL^{+}24, PW21, Shi20]. Basic [Boo21, Fle22, Knu21, Par20, Wel21, You21]. Basis [Spi23]. Basketball [JN24, YRD22]. Bayes [BM20, Bic21b, BMM21, CG23, CMB20, FG20, Gho20, Kli22, LMB21, Lig21, SDZMH24]. Bayes-Frequentist [Bic21b]. Bayesian [BS20a, Bic21a, CTÁM22, DH22b, GLW20, GZY23, GH23, HSLY23, HYW23, Har22a, HJB21, KMG^{+}21, KBSM23, KGL^{+}21, Li22, LGV23, LBM21, Mul23, PLBM24, PW21, TDM23, Hah23]. Bayesian-Frequentist [HSLY23]. Be [ZZZ21, WSCP22]. Bécue [Ben21]. Bécue-Bertaut [Ben21]. Behavior [Kva21]. Being [FH21]. Benefit [Wan23]. Benefit-Risk [Wan23]. Benjamin [Fou20]. Berger [Fou20, MH21]. Bertaut [Ben21]. Better [ZBB^{+}22]. Between [GWJ22, SH22a, Zha21a]. Beyond [Coh24, Li22, Lee21]. Bias [BdBG^{+}22, Bla20, FSOW21, LAMM22, PLBM24]. Biased [wCW24]. Biases [LD21]. Bibliographical [KP20]. Bidding [Kva21]. Big [Per22b, Qiu20]. Binary [CCL^{+}24, De 20, JSHZ21, PS21]. Binge [GBB20]. Binge-Watching [GBB20]. Binomial [And23, LR20, Ngu21, SH22a]. Biology [FA20]. Biopharmaceutical [LJ20]. Bivariate [wCW24, WA22]. Black [KBSM23, SLS22]. Blockmodels [MSS21]. Board [JNCP24]. Boca [Ben21, BS20b, Che21, CC24, Joh20, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Rot22, Sos24, Ste22, Wal24, Wel20, You21]. Böhning [MV20]. Book [Bar22, Ben21, Boo21, BS20b, Che21, Che23, CC24, Cin23, CF22, Fle22, FA20, Hah23, Joh20, Kim20b, Kim23, Knu21, Lee21, Lun20, MV20, Mar24, Mic24, Ni22, Ni24, Par20, Par23, Per22a, Per22b, Rot22, Shi20, Sos24, Ste22, Wal24, Wel20, Wel21, Wes23, You21]. Books [Boo21, Fle22, Knu21, Par20]. Bootstrap [XGH^{+}20]. both [LAMM22]. Bottai [Chr22a, Lip23]. Boundary [DP20]. Bouveyron [Shi20]. Bower [RL23]. Box [SLS22, KBSM23]. Brendan [Shi20]. Brown [Wal24]. Bryan [Rot22]. Bunge [MV20].

C [Par23, You21]. Calculating [McI22]. Calibrated [Bic21b, Bic21a]. Calibration [PLW24]. Call [BW24]. Cambridge [FA20, Shi20]. Can [EM22, Qiu20, WSCP22, ZZZ21]. Canceled [JN24]. Cannibalism [GBB20]. Cantor [Pre22]. Capture [MV20]. Care [PS21]. Carlo [Tan24, TT21]. Cary [Wel21]. Case [AG24, BHC^{+}23, Che21, FH21, GH23, RL23, WSCP22]. Case-Control [WSCP22]. Cases [MG24]. Categorical [BP23]. Cauchy [LLZL23]. Causal [AB23, CCL^{+}24, GHK24, GFB^{+}20, KS21, dMSNK23]. Causality [BS20a]. Cautionary [HY23]. Celeux [Shi20]. Censored [AG24]. Centering [GGP20]. Centers [LBV^{+}22]. Challenge [BKKL24]. Challenges [BH21, LSS^{+}24, PLBM24]. Cham [Per22a]. Change [LBV^{+}22, MCPGM23]. Change-Point [MCPGM23]. Changes [CTÁM22]. Chapman [Ben21, BS20b, Che21, CC24, Joh20, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Rot22, Sos24, Ste22, Wal24, Wel20, You21]. Characteristics [HAT24]. Characterization [VH24]. Charles [Shi20]. Charts [Fle21a]. Checking [Bic21a]. Chen [WA22]. Cheng [Knu21]. Children [PS23]. Christensen [GH22b]. Christophe [Che21]. Chun [Kim23]. City [LL20]. Claassen [Wel21]. Claims [GFB^{+}20]. Clair [Rot22]. Class [Gel23]. Class-Participation [Gel23]. Classic [KKS^{+}21]. Classical [O’N21]. Classification [LCHL22, Shi20]. Classroom [LZ21]. Clinical [Bar22, GWJ22, HAT24]. Cluster [GFHK20]. Clustering [BP23, Shi20]. Cochran [MG24]. Coefficient [HJQ20a, HJQ20b, KX20, PS21, PW21]. Coherent [HR23, Pes20]. Coin [CG23]. Collaborators [Ano20b]. Collinearity [GGP20]. collisions [Hof21]. Combination [LLZL23]. Comes [NYM24]. Comment [BLW23b, Chr20, Chr22a, Chr22b, Chr24, Fou20, GGP20, Har22b, HJQ20b, KX20, Lip23, MH21, RL23, Sch23, VW21, Tar20]. Comments [GH22b]. Committees [Bar22]. Communicating [FK21]. Community [HL22]. Community-Wide [HL22]. Companion [AB23]. Comparative [SH22b, TDM23]. Comparing [GS22, KS21]. Comparison [LGV23, Lig21]. Comparisons [BD21, Par23]. Complete [AG24]. Completely [Zha24]. Complex [KM20]. Component [LJY22]. Composite [BP23]. Compound [LZ20]. Computation [Hah23, O’N21]. Computer [FG20]. Computing [Ni24, WA22]. Concise [Sos24]. Conditional [DP20, HCS24, JN22]. Confidence [And23, BLW23a, FZ24, Lar24, LR20, PLW24, Sch23, Seg21, WMX^{+}23, WLL20]. Confirmatory [KGL^{+}21]. Conformal [JN24]. Confounding [HB22b, ZH22]. Connecting [LMB21]. Connection [GWJ22]. Connections [De 20, KP20]. Constrained [LZ20, MWM22]. Constraints [BMM21]. Consultancy [FIP^{+}23]. Consulting [LBV^{+}22]. Contestant [Kva21]. Continuous [MFD^{+}23]. Continuum [Bic21b]. Contrasting [LMB21]. Control [AB23, Qiu20, Ros22, WSCP22]. Controlled [LSS^{+}24]. Convergence [CMB20]. Cordes [Kim20b, Per22b]. Correction [Ano20a, Ano23, FSOW21]. Correlated [dSDUOG23]. Correlation [BKL^{+}22, Chr22a, GdL23, HJQ20a, HJQ20b, JSHZ21, KX20, Lip23, Wal20, dSDUOG23, PW21]. Correlation-Based [BKL^{+}22, Chr22a, Lip23]. Correlational [dMSNK23]. Correlations [Zha21a]. Count [Bla20]. Count-Data [Bla20]. Counting [HCS24]. Counts [MB20]. Course [Bil22, Cin23, Hit24]. Courses [HDM21]. Covariance [FS20]. Covariate [AG24, KS21, LAMM22]. Covariates [JN22]. Coverage [FZ24]. COVID [CTÁM22, LBV^{+}22, Li22]. COVID-19 [CTÁM22, LBV^{+}22, Li22]. CRC [Ben21, BS20b, Che21, CC24, Joh20, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Rot22, Sos24, Ste22, Wal24, Wel20, You21]. Credible [RY22]. Criteria [AMS22]. Criterion [dSDUOG23]. Cross [Wan23]. Cross-Disciplinary [Wan23]. Cui [Par23]. Curiosity [LH21]. Current [DH22b]. Curriculum [KH24]. Curve [LZ21]. Curves [vDJ20]. Customer [SCW20].

D [Fou20, Wel21]. Dankmar [MV20]. Darrin [Rot22]. Data [BP23, Ben21, Bla20, Boo21, BvdH24, De 20, EIS20, FIP^{+}23, Fle21a, GH23, HY23, JSHZ21, KBS22, KH24, Kim20b, KPC^{+}20, LR20, Lun20, Mar24, MFD^{+}23, NYM24, Ni22, PLBM24, Per22a, Qiu20, Ram23, Rot22, Shi20, SH22a, Wal24, YW20, dSDUOG23, Par20, Per22b, Bar22]. Datasets [NMVBB20]. David [Bar22, Lun20, Par20]. Day [PS21]. Decision [JNCP24, Kim20a]. Decision-Making [JNCP24]. Decision-Theoretic [Kim20a]. Decisions [Ste22]. Deconfounding [ZH22]. Deductive [dMSNK23]. Deep [PE24, Ni24]. Default [vZG22]. Defended [Bic21a]. Defense [LCHL22]. Defensive [NYM24]. Definitive [Joh20]. DeMets [Bar22]. Democratization [KPC^{+}20]. Demystifying [HDDOV22]. Dendrograms [KB23]. Density [MWM22]. Dependence [HY23, LLZL23, TWET21]. Dependent [BMM21]. Derivation [KGL^{+}21, Pre22]. Design [AB23, KMG^{+}21, LZN^{+}24, TDM23]. Designs [HAT24, KM20, KKS^{+}21, PE24]. Detecting [MGM20]. Detection [Ber21, MH20, MH21, MCPGM23, PLBM24]. Determine [NMVBB20].

Development [CF22]. Diagnostic [War23]. Diagnostics [GGP20]. Dice [AM23, Fle22]. Dickey [MWM22]. Dickhaus [Par23]. Did [Hoa20]. Difference [SH22a]. Differences [Ros22]. Different [GHK24]. Differentiable [HGK^{+}22]. Differentially [YR24]. Dimensional [BvdH24, JSHZ21]. Dimensions [Tan24]. Ding [Par23]. Directional [Ber21, Che21, MH20, MH21]. Directionality [MGM20]. Disciplinary [Wan23]. Disciplines [HSHM^{+}23]. Discovery [HL22]. Discrepancy [MFD^{+}23]. Discrete [Hua23]. Discrimination [MG24]. Discussion [Hit24]. Discussion-Based [Hit24]. Diseases [Li22]. Display [Hoa20, Ros22]. Dissertations [FIP^{+}23]. Distributed [Cha23]. Distribution [Arn20, wCW24, DP20, HB22a, KP20, LD21, O’N21, PGW22, Pre22, SKH23]. Distribution-Free [SKH23]. Distributional [RKK24]. Distributions [Hua23, Lar24, Lig21, NR22]. Distributive [Rei21]. divergence [PW21]. divergence-based [PW21]. Do [FZ24, Gra23, Fle22]. Does [JDC20, PS21]. Don’t [Gra23]. Doomed [CZ22]. Drop [KMG^{+}21]. Drop-the-Losers [KMG^{+}21]. Drug [CF22, Wan23]. During [CTÁM22].

e-book [Ben21, Lee21, You21]. Early [Fre21]. Earthquake [AW24]. ed [Bar22, CF22, Fle21b, Mic24, Ni22, Per22a, Sos24]. ed. [Che23, Cin23]. Editor [KX20, GWJ22, Hoa20, HJQ20b, MH21, VW21, Zha21c]. Editorial [Ano20b]. Editors [Ano24]. eds [MV20, Par23]. Education [DH22b, KPC^{+}20, RUCH23]. Effect [CCZ24, CCL^{+}24, GHK24, KBF21]. Effects [Ber21, LBM21, MH20, MH21, Mul23]. Efficient [HDDOV22, JSHZ21]. Eigenvalues [Wal20]. EIV [LZ20]. Elizabeth [Wel21]. Ellenberg [Bar22]. Emphasis [Mar24]. Empirical [KS21, XC23]. Endogeneity [NCK^{+}22]. Engineering [BS20b]. Enhanced [GZWL24]. Entity [BYH^{+}23]. Entropy [VW21, Zha21b]. Envelopes [War23]. Environment [HHL21]. Epidemic [CTÁM22]. Equality [FS20, MWM22]. Equations [HB22b]. Equivalence [SH22b]. Erroneous [GFB^{+}20]. Error [VW21, Zha21b, vZG22]. Errors [Bic21b, GZWL24, LAMM22]. Estimate [Cha23]. Estimates [CCZ24, KBF21, LR22, LD21, vZG22]. Estimating [Bic21b, BYH^{+}23, HB22b, MKR^{+}23, PGW22]. Estimation [CG23, CCL^{+}24, DP20, GZWL24, HWM24, HJQ20a, HJQ20b, HCS24, KX20, SDZMH24, VW21, Zha21b, Zha24]. Estimator [BM20, GFY22, QSL23]. Estimators [Ano23, GGK20, LJY22, Por22, SDZMH24]. Ethical [FH21]. Ethics [RUCH23]. Eugenia [Knu21]. Evaluating [YRD22]. Evaluation [Wan23]. Event [Che23]. Events [LL20]. Evidence [Pes20, Spi23]. Evidential [PLW24]. Exact [Lig21, RP21]. Example [And23, Gho20]. Examples [Arn20, Che20, Zha21a, Mic24]. Exceedance [Seg21]. Excellence [Per22a]. Expectancy [PG23]. Experiment [Gra20]. Experiments [HDM21, JN22, KKS^{+}21, LSS^{+}24, Zha24]. Expert [FH21]. Explained [ZZZ21]. Exploration [KB23]. Exploratory [LD21, Ni22]. exploring [SRSK22]. Expressing [RY22]. Extension [SK20]. Extensions [KP20, Oga20]. Extreme [EIS20, LL20].

Faces [Che24]. Facilitating [Fre21]. Fact [Har22a]. Factor [FG20, LMB21, RP21]. Factors [BMM21, CG23, CMB20, Kli22, Lig21]. Fail [CZ22]. False [HL22]. Fast [BLW23a, BMS21, Sch23]. Feature [BS20b]. Felix [Sos24]. Few [EIS20]. Fiction [Har22a]. Fields [BMS21]. Finance [Mic24]. Finite [GZWL24]. Fires [AGK^{+}21]. First [CTÁM22, Cin23, Coh24]. First-Passage [Coh24]. Five [KS21]. FL [Ben21, BS20b, Che21, CC24, Joh20, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Rot22, Sos24, Ste22, Wal24, Wel20, You21]. Fleming [Bar22]. Flexible [HGK^{+}22]. Follmann [McI22]. Football [BW24, NYM24]. Forbidden [Chr24, Roh23]. Forecast [BKKL24]. Forecasting [BKKL24]. Forecasts [YRD22]. Forest [SLS22, ZZNN20]. Forget [Han20]. Form [RP21]. Fostering [YW20]. Foulley [Fou20]. Found [GWJ22]. Fractional [XGH^{+}20]. Fractional-Random-Weight [XGH^{+}20]. Frailty [Cha23]. Framework [NMVBB20]. Frameworks [LZ20]. Frans [Mar24]. Free [SDZMH24, SKH23]. Frequency [AW24, vDJ20]. Frequentist [Bic21b, HSLY23]. Fresh [Rot22]. Function [Cha23, QSL23]. Functional [MKR^{+}23]. Functions [wCW24, FK21, HDDOV22, SW21]. Fundamentals [Fle21b]. Further [Arn20, Zha21a]. Future [DH22b].

G [MV20, Sch23]. GAISE [RUCH23]. Game [Gra20, SRSK22]. Garrett [Joh20]. Gary [Kim20b, Per22b]. Gaussian [BMS21, KP20, Mul23]. Gender [Knu21]. General [CMB20]. Generalization [Muk22, MWM22]. Generalize [CCZ24]. Generalized [HB22b, Lee21]. Generate [JSHZ21]. Generating [NR22, SW21, Wal20]. Generation [ZSB22]. generators [Hof21]. Gengsheng [KX20]. Geometric [PGW22, Pre22]. George [Wel21]. Geostatistical [De 20]. Ghahramani [Fle21b]. Gilles [Shi20]. Global [War23]. Glucose [MFD^{+}23]. God [Fle22]. Going [GBB20]. Golbeck [Per22a]. Government [FSOW21]. Graduate [Bil22, JDC20]. Grant [Fle21a]. Graph [Kim23]. Graphical [Ros22, vDJ20]. Graphics [Fle21a]. Griffiths [Chr22b, Har22b]. Grolemund [Joh20]. Groups [GS22]. Guaranteed [FZ24]. Guidance [HSHM^{+}23]. Guide [Joh20]. Guidelines [RUCH23].

H [Bar22, Ben21, BS20b, CC24, CF22, Fle22, Kim20b, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par20, Par23, Per22a, Per22b, Rot22, Shi20, Sos24, Ste22, Wal24, You21]. H/e [Lee21, You21]. H/e-book [Lee21, You21]. Habiger [Ber21]. Haenszel [MG24]. Hall [Ben21, BS20b, Che21, CC24, Joh20, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Rot22, Sos24, Ste22, Wal24, Wel20, You21]. Hall/CRC [Ben21, BS20b, Che21, CC24, Joh20, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Rot22, Sos24, Ste22, Wal24, Wel20, You21]. Hamiltonian [TT21]. Handbook [Par23]. hardback [Boo21, Che21, Knu21]. Harmonizing [KKS^{+}21]. Harville [GH22b]. Hazard [FO21]. Heads [FSOW21]. Health [NCK^{+}22, PS21]. Heijden [MV20]. Help [Qiu20]. Henk [Wel20]. Here [NYM24]. Heterogeneity [CTÁM22]. Hidden [AW24, GFB^{+}20]. Hierarchical [MCPGM23]. High [BvdH24, JSHZ21, Tan24]. High-Dimensional [BvdH24, JSHZ21]. Hill [Chr22b, Har22b]. Historical [PLBM24, vDJ20]. History [Che23, HB22a, Hit24]. Hitting [AM23, Che24]. Hoboken [Bar22, CF22]. Holmes [FA20]. Home [PS21]. Hospital [PS21]. Houston [EM22]. Hsu [Par23]. Hu [KX20]. Huber [FA20]. Human [CTÁM22]. Hurdle [MSS21]. Hybrid [HSLY23]. Hypergeometric [BLW23a, Sch23]. Hypotheses [Chr22b, GH22a, GH22b, HR23, Har22b, LMB21, MWM22]. Hypothesis [Bic21b, Bic21a, GOIRS21, Kim20a, MFD^{+}23].

Ian [Fle22]. Identical [Arn20]. Identities [WA22]. Identity [SK20]. If [ZZZ21]. Illuminate [SLS22]. Illuminating [And23]. Illustrating [HDM21]. Illustrations [Muk22]. Imbalances [JN22]. Impact [Ano22, SLS22, YK21]. Implication [LH21]. Implications [AG24, NCK^{+}22]. Imply [Muk22]. Importance [LZN^{+}24]. Improved [GdL23]. Improvements [Oga20]. Improving [BB19, Fou20, KBF21]. Imputation [BvdH24, Che20, LR22, LR20, NMVBB20, SH22a]. Imputed [NMVBB20]. Inadmissible [BM20]. Inclusive [Per22a]. Incomplete [SH22a]. Inconsistency [FO21]. Incorporating [AMS22]. Increase [PS21]. Increased [FSOW21]. Independence [Muk22]. Index [DND^{+}20]. Induced [NCK^{+}22]. Industry [LJ20]. Inequities [NCK^{+}22]. Infectious [Li22]. Inference [AB23, Bla20, GZWL24, HSLY23, Har22a, HWM24, HJB21, JN22, KBS22, KS21, KH24, LBM21, LR22, LJY22, You21, CC24]. Inferences [GFB^{+}20, HHL21]. Infinitely [Che24]. Inflated [HCS24]. Influence [FK21, HDDOV22]. Information [CCL^{+}24]. Informative [vZG22]. Informed [GLW20]. Inspection [RK23]. Institute [Wel21]. Instruction [RUCH23]. Integer [LR22]. Integer-Valued [LR22]. Integrated [CCL^{+}24]. Integrating [RUCH23]. Integration [Tan24]. Interactive [Fle21a, FG20, KB23]. Intercept [Cha23]. Interpreted [Bic21b]. Interpreting [PS21]. Interval [And23, HR23, HJQ20a, HJQ20b, KX20, LMB21]. Intervals [BLW23a, FZ24, LR20, PLW24, RY22, SLS22, Sch23, Seg21, WMX^{+}23, ZZNN20]. Introducing [KH24]. Introduction [Wal24, Wel21, Wes23, Sos24, CC24, Mic24]. Intuition [FK21, LH21]. Invariant [VW21, Zha21b]. Inverse [KBF21, Zha24]. Investigation [EM22]. ISBN [Bar22, Ben21, Boo21, BS20b, Che21, CC24, CF22, Fle22, FA20, Joh20, Kim20b, Kim23, Knu21, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par20, Par23, Per22a, Per22b, Rot22, Shi20, Sos24, Ste22, Wal24, Wel20, Wel21, You21]. Issues [CF22]. Item [SCW20]. Iterative [NMVBB20]. IVEware [Che20].

J [Fou20, Joh20, Sch23, VW21]. J. [Zha21c]. Jackknife [Ano22, YK21]. Jason [Par23]. Jay [Kim20b, Per22b]. Jeffrey [Ni22]. Jeffreys [Gho20].

Jeremy [You21]. John [CF22, MV20, JDC20]. Johnson [BS20b, vDJ20]. Joint [LZN^{+}24]. Jolani [Mar24]. Journals [HSHM^{+}23]. Julie [Lee21]. Jung [KX20]. Justice [Rei21].

Kaplan [GFY22, QSL23]. Kendall [vdHZ22]. Keydana [Ni24]. Kingdom [Kim20b, Per22b]. Kjell [BS20b]. Knee [MKR^{+}23]. Know [Boo21, Gra23]. Knowledge [Chr24, Roh23]. Krishnamoorthy [HJQ20b]. Kuhn [BS20b].

L [Bar22, Ni22, Per22a, Sch23]. Lambert [Cha23]. Laplace [KP20]. Large [KB23, LBM21]. Lasso [Rei21]. Leadership [Per22a]. Leaf [Hoa20]. Lean [BBB^{+}21]. Learn [Par20]. Learned [BYH^{+}23]. Learning [BdBG^{+}22, BKKL24, HDDOV22, KS21, Ni24, PE24, SW21, TDM23, TT21]. Least [GGK20]. Least-Square [GGK20]. Led [GFB^{+}20]. Legal [FH21]. Legler [Lee21]. Lens [VPH23]. Lessons [BYH^{+}23, Hit24]. Lest [Han20]. Letter [GWJ22, Hoa20, HJQ20b, MH21, VW21, Zha21c]. Letters [KX20]. Level [LL20]. Ley [Che21]. Li [Kim23]. Li-Chun [Kim23]. Lie [Gel23]. Life [PG23]. Likelihood [BP23, HCS24, MG24, SDZMH24, WLL20]. Likelihood-Free [SDZMH24]. Likelihood-Ratio-Based [WLL20]. Limit [QSL23]. Limiting [KBF21, vDJ20]. Linear [Ano23, Chr22b, Chr24, Cin23, GH22a, GH22b, Har22b, Lee21, Mar24, Mul23, Por22, Roh23, SLS22, ZH22, vdHZ22]. Linearity [Por22, Ano23]. Linked [KBS22]. Listings [GH23]. Location [SKH23]. Location-Scale [SKH23]. Log [GZY23]. Log-Rank [GZY23]. Logistic [PS21]. Long [FO21]. Long-Term [FO21]. Longford [Ste22]. Longitudinal [Mar24]. Look [VPH23]. Looking [EIS20]. Lorden [Sch23]. Lorenz [LZ21]. Losers [KMG^{+}21]. Loss [GGK20, PG23]. Lotteries [Han20, JDC20]. Low [AW24]. Low-Frequency [AW24]. Ltd [CF22]. Luminaries [JDC20]. Lumley [BC23].

M [MV20]. Machine [BdBG^{+}22, KS21]. Madness [ZZZ21]. Main [Chr24, Roh23]. Make [Boo21]. Making [HJB21, JNCP24, Ste22]. Man [Gal20]. Mandel [Tar20]. Manifesto [Knu21]. Mantel [MG24]. Many [Che24, HHL21, Hof21]. Mapping [PG23]. Maps [Fle21a]. March [ZZZ21]. Markdown [Joh20]. Market [Mic24]. Markov [AW24, BMS21, Oga20]. Martinez [Ni22]. Mass [QSL23]. Matched [MFD^{+}23, Ros22, WSCP22]. Matched-Pair [Ros22]. Matching [KS21]. Mathematical [LJ20]. Mathematician [Knu21]. Mathematics [Fle22]. MATLAB [Ni22]. Matrices [FS20, Wal20]. Matrix [GdL23]. Matters [GWJ22, Van21]. Mavrakakis [You21]. Max [BS20b]. Maximum [MFD^{+}23, SK20]. May [Muk22]. McCann [Ber21]. MCMC [LH21]. Mean [Gho20, MFD^{+}23, VW21, Zha21b]. Measure [VW21, Zha21b]. Measures [Ber21, MH20, MH21, Pes20]. Median [Ano22, YK21]. Medical [MV20]. Medicine [DH22a, GFB^{+}20]. Meier [GFY22, QSL23]. Melded [FZ24]. Memoryless [NR22]. Method [LJY22, SH22a, Wal20]. Methodology [LSS^{+}24]. Methods [Che21, JSHZ21, KS21, LCHL22, MV20, YR24]. Metrics [SH22b]. Micha [Tar20]. Milliken [Wel21]. Miltiadis [You21]. Minimum [AMS22]. Minimums [SK20]. Mirage [Per22b]. Misclassification [GZWL24, LAMM22]. Missing [BvdH24, EIS20, LR20, Mar24, MFD^{+}23]. Mixed [Wel21, Wes23]. Mixture [BP23, KP20, LJY22]. Mobility [CTÁM22]. Model [Ano23, Bic21a, Boo21, BD21, CCL^{+}24, Chr22b, Cin23, Coh24, FO21, GH22a, GH22b, Har22b, KBSM23, LJY22, MB20, Man20, Por22, RP21, Tar20, Shi20]. Model-Based [Shi20]. Modeling [AGK^{+}21, BH21, BW24, FG20, HGK^{+}22, Hah23, MSS21, NCK^{+}22, SCW20]. Modelling [CC24]. Models [AW24, AG24, Bla20, Cha23, De 20, HCS24, Lee21, LZ20, LGV23, MKR^{+}23, War23, WLL20, Wel21, Wes23, ZH22, dMSNK23, BS20b, You21]. Modern [Che21, FA20, FA20]. Modified [LMB21]. Moments [LJY22, Ngu21, WA22]. Mónica [Ben21]. Monitoring [Bar22, MFD^{+}23]. Monotonic [vdHZ22]. Monte [TT21, Tan24]. Mortality [FSOW21]. Most [VH24]. Movement [MKR^{+}23]. MOVER [WMX^{+}23]. MOVER-R [WMX^{+}23]. Multi [HPZ23]. Multi-Response [HPZ23]. Multilevel [Lee21, MKR^{+}23]. Multinomial [Chr20, Lig21, Sza19]. Multiple [Che20, CCL^{+}24, Lee21, LR22, LR20, NMVBB20, Par23, SH22a]. Multiple-Model-based [CCL^{+}24]. Multivariate [LGV23, McI22]. Murphy [Shi20]. Murray [CC24]. My [Swa20]. Myth [MWS23]. Myths [vdHZ22].

Name [PS23]. National [YRD22]. NC [Wel21]. NCAA [JN24]. Need [Boo21]. Network [MSS21]. Networks [BP23]. Neural [SDZMH24]. Neutralizing [DND^{+}20]. Newcombe [SH22a]. Nicholas [Ste22]. Nicolas [Mic24]. NJ [Bar22, CF22]. Nonlinear [Mul23]. Nonnegligible [Ber21, MH20, MH21]. Nonparametric [DP20, LZ20, VH24]. Nonparametrics [BHC^{+}23, RL23]. Normal [Gal20, Gho20, LGV23, LD21]. Normally [Cha23]. Note [CMB20, GGP20, SCW20, Tan24]. Null [Bic21b, Bic21a, HR23, LMB21]. Number [LBM21, NMVBB20, Hof21]. NY [Fle22, Par20].

O [Fou20]. Object [Ram23]. Objective [PW21]. Observational [LZN^{+}24, WSCP22]. Observations [Mar24]. Observed [JN22]. Occupancy [O’N21]. Ockham [RP21]. Old [Gra20, LH21]. One [BD21, CCZ24, HHL21, McI22, SY21]. One- [SY21]. One-Sided [McI22]. One-Step [CCZ24]. One-Way [BD21]. Online [LSS^{+}24, SCW20]. Operating [HAT24]. Optimal [BLW23a, BKL^{+}22, Chr22a, Lip23, Sch23]. Optimality [AMS22]. Optimized [KKS^{+}21]. Optimizing [KMG^{+}21]. OptSig [Kim20a]. Order [BMM21, MWM22]. Oriented [Ram23]. Out-of-Sample [HWM24]. Outcome [CCL^{+}24, Chr20, Sza19, WSCP22, YRD22]. Outcomes [BW24, BMM21, CCL^{+}24]. Overdispersed [MB20]. Overfitting [Chr24, Roh23]. Oxford [Kim20b, Per22b].

P [FA20, Joh20, Wel20]. Package [Kim20a]. Page [Boo21]. Pair [Gra23, Ros22, WSCP22]. Paired [Gra23, HY23]. Pairs [MFD^{+}23]. Pairwise [BD21, Muk22]. Pandemic [LBV^{+}22]. Paper [Swa20]. paperback [Wel21]. Paradox [Ald21, RK23, DND^{+}20]. Parallel [LBM21]. Parameter [HYW23, Lar24, PGW22, SDZMH24, WLL20]. Parameterization [FO21]. Parameterizations [GFHK20]. Parity [BdBG^{+}22]. Parole [JNCP24]. Partial [Coh24]. Participation [Gel23]. Pass [NYM24]. Passage [Coh24]. PatentsView.org [BYH^{+}23]. Pattern [Per22b]. Patterns [MKR^{+}23, SW21]. Paul [Lee21]. Peak [MWS23]. Pearson [vdHZ22]. Penalized [BP23, WMX^{+}23]. Penney [SRSK22]. Penney-Ante [SRSK22]. Penzer [You21]. Percentage [GWJ22, Van21]. Peremptory [PLBM24]. Perform [Gra23]. Performance [BYH^{+}23, DND^{+}20]. Permutation [CZ22]. Personalized [GFB^{+}20]. Perspective [Bar22]. Perspectives [KGL^{+}21, vDJ20]. Peter [MV20]. Phantom [Per22b]. Phistogram [Bla24]. Phlegon [Hoa20]. Pitfalls [Kim20b]. Planning [GFHK20, Per22a]. Play [BW24, Fle22, XC23]. Player [NYM24]. Plot [LZN^{+}24, PE24]. Plots [AB23, War23]. Plotting [WLL20]. Point [LR22, MCPGM23]. Points [DP20]. Poisson [HB22a, WA22]. Policies [HL22]. Pool [Spi23]. Population [CCZ24, GZWL24]. Positive [TWET21]. Posterior [CG23, LBM21, SY21, SH22b]. Power [Chr22b, GH22a, GH22b, HYW23, Har22b, KMG^{+}21]. Powerful [VH24, WSCP22]. pp [Bar22, Ben21, Boo21, BS20b, Che21, CC24, CF22, Fle22, FA20, Joh20, Kim20b, Kim23, Knu21, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par20, Par23, Per22a, Per22b, Rot22, Shi20, Sos24, Ste22, Wal24, Wel20, Wel21, You21]. Practical [Bar22, BS20b, KBS22, SH22b, Wan23]. Practice [Che20, Fre21, LBV^{+}22]. Pragmatic [LJ20]. Predict [JN24]. Prediction [Ald21, BKL^{+}22, Chr22a, LGV23, Lip23, SLS22, YRD22, ZZNN20]. Predictions [HPZ23]. Predictive [BS20b]. Preferential [VPH23]. Presence [GOIRS21]. Preservation [GH23]. Press [Ben21, BS20b, Che21, CC24, FA20, Joh20, Kim20b, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Per22b, Rot22, Shi20, Sos24, Ste22, Wal24, Wel20, You21]. Prevalence [GZWL24]. Price [Kva21]. Prime [AM23, Che24]. Primer [Kim20a]. Principles [You21]. Prior [HYW23]. Prioritization [KS21]. Prioritizing [LZN^{+}24]. Priors [Kli22, Mul23, vZG22, PW21]. Prism [BdBG^{+}22]. Privacy [GH23]. Private [YR24]. Privault [Mic24]. Probabilistic [BKKL24, HPZ23, Ngu21, YRD22]. Probabilities [Bic21b, SH22b]. Probability [Fle21b, JN24, PS21, QSL23, Seg21, SY21, SH22b, WLL20, Wel20, Zha24, Rot22, You21]. Problem [HHL21, PS23, Per22b, Rei21, VPH23, dMSNK23]. Procedure [LMB21]. Process [Mul23, Qiu20]. Processes [Fle21b]. produce [Hof21]. Product [QSL23]. Product-Limit [QSL23]. Profiles [EIS20]. Programming [HGK^{+}22]. Projections [Wal20]. Propensity [KBF21]. Proper [BM20, KBS22]. Properties [De 20]. Property [NR22]. Proportions [HCS24, LR20, SH22a]. Proposal [vZG22]. Protection [GH23]. Prototypes [KB23]. Provide [FZ24]. Pseudo [ZBB^{+}22]. Pseudo-Ranks [ZBB^{+}22]. Publication [HL22]. Publish [Swa20]. Published [LD21]. Python [Wal24, Hah23].

Qin [KX20]. Quadrant [TWET21]. Quantifying [RK23]. Quantitative [Wan23]. Quantities [WMX^{+}23]. Quartets [GHK24]. Quetelet [Gal20].

R [Bar22, Bus21, Ni22, Ni24, Wal24, Kim20a, TT21, WMX^{+}23, Ben21, Che23, Lee21, Lun20, Rot22, Shi20, Joh20, Wes23]. RafterNet [HPZ23]. Raftery [Shi20]. Raise [JDC20]. Random [BMS21, Cha23, Coh24, Hof21, Ngu21, RK23, SLS22, SW21, XGH^{+}20, ZZNN20]. Randomization [KKS^{+}21, Hem24]. Randomized [GFHK20, Zha24]. Randomness [HDM21]. Rank [GZY23]. Ranked [HSHM^{+}23]. Ranking [ZBB^{+}22]. Ranks [BD21, Zha21a, ZBB^{+}22]. Rare [WSCP22]. Rate [HL22]. Ratings [SCW20]. Ratio [FO21, MG24, MWM22, WMX^{+}23, WLL20]. Raton [Ben21, BS20b, Che21, CC24, Joh20, Kim23, Lee21, Lun20, MV20, Mar24, Mic24, Ni24, Par23, Rot22, Sos24, Ste22, Wal24, Wel20, You21]. Re [SRSK22]. Re-exploring [SRSK22]. Readmissions [PS21]. Real [YRD22, ZZZ21]. Real-Time [YRD22]. Recapture [MV20]. Recommendations [BB19, DH22b, Fou20]. Reconnecting [SY21]. Reconstructing [GFY22]. Recreational [MKR^{+}23]. Recruitment [MWS23]. Recurrence [AW24]. Recursive [HJB21]. Reduction [Spi23]. Reenactments [HB22a]. Regions [WLL20]. Regression [AG24, BBB^{+}21, Chr24, DP20, HPZ23, Lee21, LZ20, LGV23, Mar24, MKR^{+}23, PS21, Roh23, RKK24, SLS22, SKH23, War23]. Regret [RY22]. Reinforcement [TDM23]. Rejoinder [GH22b]. Related [Arn20, Zha21a]. Relative [DND^{+}20, VW21, Zha21b]. Releasing [YR24]. Repaired [JDC20]. Replicability [HHL21, Seg21]. Replicable [HL22]. Reporting [Spi23]. Representation [KP20]. Reproducibility [XC23]. Research [HL22]. Residential [NCK^{+}22]. Resolution [BYH^{+}23]. Response [AGK^{+}21, BLW23b, BC23, HPZ23, HJQ20b, LAMM22, MH21, SCW20, Zha21c]. Restricted [Kli22]. Results [YR24]. Rethinking [Knu21]. Return [LL20]. Reversals [PS21]. Review [Bar22, Ben21, Boo21, BS20b, Che21, Che23, CC24, Cin23, CF22, Fle22, FA20, Hah23, Joh20, Kim20b, Kim23, Knu21, KGL^{+}21, LSS^{+}24, Lee21, LCHL22, Lun20, MV20, Mar24, Mic24, Ni22, Ni24, Par20, Par23, Per22a, Per22b, Rot22, Shi20, Sos24, Ste22, Wal24, Wel20, Wel21, Wes23, You21]. Revisited [KP20, Man20, Tar20]. Revisiting [Gho20, PS23]. Rice [BC23]. Right [AG24, Kva21]. Right-Censored [AG24]. Risk [JNCP24, Wan23]. Ritov [Sos24]. Roback [Lee21]. Robert [Fle21a]. Robust [CCL^{+}24]. Robustified [LBM21]. Robustness [GGK20]. Roger [MH21]. Rohlfs [Chr24]. Rolling [Che24]. Rolls [AM23]. ROPE [LMB21]. Row [Kva21]. Runners [MKR^{+}23]. Rush [NYM24]. Russell [Wel21].

S [Bar22, Lun20, Mar24]. Saeed [Fle21b]. Safety [Wan23]. Same [CG23, GHK24]. Sample [Ano22, Bla20, HSLY23, HWM24, KMG^{+}21, KGL^{+}21, McI22, YK21]. Samples [EIS20]. Sampling [BMS21, wCW24, GZWL24, VPH23, Kim23]. Sampling-Based [GZWL24]. SAS [Wel21, Wel21]. Savage [MWM22]. Scale [SKH23]. Scaled [Man20, Tar20, vZG22]. Scenarios [HAT24]. Schilling [BLW23b, Sch23]. Science [Ben21, FIP^{+}23, KH24, KPC^{+}20, Per22a, Shi20, YW20, Kim20b]. Sciences [MV20]. Scientific [HSHM^{+}23, Ni24]. Score [KBF21]. Scott [Boo21]. Second [ZSB22]. Second-Generation [ZSB22]. Secondary [CCL^{+}24]. See [EIS20]. Segregation [NCK^{+}22]. Selecting [HAT24]. Selection [BS20b, Har22a, RP21, ZSB22, dSDUOG23]. Selective [Han20, JDC20]. Semi [LJY22, RKK24]. Semi-Structured [RKK24]. Semi-Supervised [LJY22]. Seminar [Bil22]. Senn [CF22]. Sensitivity [HAT24]. Sequential [TDM23]. Series [BKKL24, Lun20, MB20, MGM20]. Service [Han20, JDC20]. Sessions [MKR^{+}23]. Set [CMB20, JSHZ21]. Set-Up [CMB20]. Sets [Kli22]. Seventeen [Wel20]. Shahab [Mar24]. Shared [Cha23]. Shining [SLS22]. Shiny [Gra20]. Short [CMB20, FO21, Wel20, Wes23]. Short-Term [FO21]. Should [Swa20]. Shumway [Lun20]. Side [Wal24]. Side-by-Side [Wal24]. Sided [McI22, SY21]. Sides [CG23]. Sign [Bic21b, HY23, PS21]. Significance [Ber21, Bic21b, Bic21a, MH20, MH21]. Sigrid [Ni24]. Simple [McI22, VH24]. Simpson [DND^{+}20]. Simulation [War23]. Size [KMG^{+}21, KGL^{+}21, McI22]. Sizes [HSLY23]. Skills [LJ20]. Slowly [EM22]. Slugging [Van21, GWJ22]. Small [Bla20, HSLY23, NCK^{+}22]. Small-Area [NCK^{+}22]. Small-Sample [Bla20]. Smith [Kim20b, Per22b]. Snow [LL20]. Social [MV20]. Solka [Ni22]. Solution [Chr24, Roh23]. Solved [Rei21]. Solving [dMSNK23]. Some [Oga20]. Sons [CF22]. Sophomore [MWS23]. Sources [Chr24, Roh23]. Sparse [MSS21]. Spatial [CTÁM22, HDM21, HB22b, ZH22]. Spatially [dSDUOG23]. Spatio [Bus21, MCPGM23]. Spatio-Temporal [Bus21, MCPGM23]. Spearman [vdHZ22]. Specialized [Chr24, Roh23]. Specified [JSHZ21, Wal20]. Speegle [Rot22]. Spiegelhalter [Par20]. Sports [Swa20]. Spotlight [SLS22]. Springer [Per22a]. Square [GGK20]. Stability [YR24]. Stage [KMG^{+}21]. Standard [vZG22]. State [DH22b, XC23]. States [CTÁM22]. Statistical [BH21, Ber21, BdBG^{+}22, CC24, FH21, HGK^{+}22, HSHM^{+}23, HHL21, HDDOV22, LSS^{+}24, LBV^{+}22, LJ20, LJY22, MH20, MH21, NCK^{+}22, Qiu20, Spi23, You21, CF22, Sos24]. Statistician [Ano24, HJQ20b, JDC20, Fou20, KX20, MH21, Sch23, Tar20, VW21]. Statistics [Bil22, Bus21, Che21, DH22a, FIP^{+}23, FA20, Fre21, HDM21, Hit24, JDC20, KH24, Par20, Per22a, RUCH23, Rot22, VPH23, VH24, XC23, Ste22]. Stein [WA22]. Stem [Hoa20]. Stem-and-Leaf [Hoa20]. Step [CCZ24]. Stephen [CF22]. Stewart [Fle22]. Stochastic [Fle21b, Mic24]. Stoffer [Lun20]. Stories [Wel20]. STRAIN [NYM24]. Strategies [BW24, BMS21]. Strategy [KM20]. Strength [Kli22, Spi23]. Strike [PLBM24]. Stroup [Wel21]. Structure [AGK^{+}21, LBV^{+}22, dSDUOG23]. Structured [RKK24]. Structures [JSHZ21, LLZL23, SW21]. Student [JDC20]. Students [Bil22, Fre21]. Studies [WSCP22, Che21]. Study [AB23, FSOW21, GH23, HYW23, LZN^{+}24, PGW22]. Style [FIP^{+}23]. Successes [BLW23a, Sch23]. Summarizing [HAT24]. Sums [Coh24]. Supervised [LJY22]. Sure [CMB20]. Surfaces [FG20]. Surprises [Wel20]. Surrogate [FG20]. Survey [BdBG^{+}22, VPH23]. Survival [FSOW21]. Susan [Bar22, FA20]. Susceptibility [HCS24]. Switzerland [Per22a]. Synthesis [GH23]. System [vDJ20]. Szabo [Chr20].

T [Shi20, Ste22]. Tailed [Pes20]. Tale [HY23]. Tan [Mar24]. Target [BW24, CCZ24]. Taylor [Wal24]. Teaching [FK21]. Team [JDC20]. Technical [PE24]. Temporal [Bus21, MCPGM23, SW21]. Term [FO21, Hem24]. Test [Chr20, CZ22, Gho20, Gra23, GZY23, Hem24, HY23, LLZL23, McI22, MG24, Sza19, VH24, Chr22b, GH22a, GH22b, Har22b]. Testing [Bic21b, Bic21a, BMM21, GOIRS21, Kim20a, LSS^{+}24, LMB21, MFD^{+}23, MWM22, Mul23, TWET21, PW21]. Tests [FS20, GLW20, HR23, SY21, SH22b]. Textual [Ben21]. Their [PS21]. Theorem [Oga20]. Theoretic [Kim20a]. Theory [Sos24, SCW20, Cin23]. Thinker [Boo21]. Thomas [Bar22, Che21]. Thorsten [Par23]. Three [BB19, Fou20, GS22]. Tijms [Wel20]. Time [AGK^{+}21, BKKL24, GFHK20, JDC20, MB20, MGM20, RK23, YRD22, Lun20]. Time-Series [MB20]. Times [Coh24]. Top [HSHM^{+}23]. Top-Ranked [HSHM^{+}23]. torch [Ni24]. Tournament [Ald21]. Tournaments [JN24]. Tracking [NYM24]. Training [Chr24, MKR^{+}23, Roh23]. Transformation [Ros22]. Transport [CCZ24]. Treated [Ros22]. Treated-Control [Ros22]. Treatment [CCZ24, CCL^{+}24, GHK24]. Trend [Chr20, LL20, Sza19]. Trial [GFHK20, HAT24]. Trials [Bar22, GWJ22, KGL^{+}21]. Truncated [LD21, PGW22]. Truths [Gel23]. Tukey [JDC20]. Tutorial [TDM23]. Two [CG23, Chr24, KMG^{+}21, LJY22, PS23, Pes20, Roh23, SY21, WMX^{+}23, WLL20, Gel23]. Two-Children [PS23]. Two-Component [LJY22]. Two-Parameter [WLL20]. Two-Sided [SY21]. Two-Stage [KMG^{+}21]. Two-Tailed [Pes20].

U.S. [Han20]. UK [FA20, Shi20]. Unaffected [Har22a]. Unbiased [Ano23, Por22]. Uncertainty [Fle22]. Underdispersed [Hua23, MB20]. Undergraduate [DH22b, Fre21, YW20]. Understanding [AG24]. Unified [RY22]. Uniform [Man20, Tar20]. United [Kim20b, Per22b, CTÁM22]. Univariate [LGV23, WLL20]. University [FA20, Kim20b, Per22b, Shi20].

Unseen [HCS24]. Update [Gra20]. Updating [BMS21]. Use [BB19, Fou20, Hoa20, PE24]. Using [BD21, Cha23, Che20, FG20, HGK^{+}22, HCS24, JN24, KS21, LD21, Lun20, MKR^{+}23, Mul23, PLBM24, Rot22, SLS22, SH22a, SH22b, Wal20, WA22, Wes23, JNCP24, LZN^{+}24]. Utility [GH23].

v [Kim20b]. Validation [PE24]. Value [GFB^{+}20, SY21]. Valued [LR22]. Values [BB19, Fou20, Kli22, Pes20, ZSB22]. Variability [KBF21]. Variable [LZN^{+}24, LAMM22, ZSB22]. Variables [LZN^{+}24, Ngu21, Zha21a]. Variances [AMS22, BMM21]. Variant [PS23]. Variational [KBSM23, KH24]. Velilla [GGP20]. Verdebout [Che21]. Versatile [Chr24, Roh23]. Versus [LGV23, Mul23]. Very [CG23]. via [KS21]. Vietnam [JDC20]. View [RY22]. vii [Per22b]. Viral [GBB20]. Visual [AB23]. Visualization [Fle21a, FG20, GdL23]. Visualizing [FS20]. Visually [FK21].

W [JDC20, Wel21]. Wald [And23]. Walter [Wel21]. Wang [Sch23]. Was [Gal20]. Watching [GBB20]. Wave [CTÁM22]. Way [BD21, ZBB^{+}22]. Ways [GHK24]. Weight [XGH^{+}20]. Weighted [AMS22]. Weighting [CCZ24, Zha24]. Weights [KBF21]. Wendy [Ni22]. Where [Swa20]. Wide [HL22]. Wiley [Bar22, CF22]. Wilson [LR20]. Win [JN24]. Winners [JN24]. Wolfgang [FA20]. Wolfinger [Wel21]. Work [Boo21]. Working [dSDUOG23]. Wu [VW21].

x [Knu21, Mic24]. Xia [HJQ20b, KX20]. Xie [Joh20]. xii [Lun20, Rot22, Wel21]. xiii [Boo21, Sos24, Wel20]. Xinjie [KX20]. Xinping [Par23]. xiv [Par23]. xix [Ni24, Wal24, You21]. xv [BS20b, Kim23, Ste22]. xvi [Che21, CC24, Par20]. xvii [Ben21, Lee21, Shi20]. xxi [CF22, Mar24]. xxii [Bar22]. xxiii [FA20]. xxvii [Per22a]. xxxii [MV20]. xxxiv [Joh20].

Ya [Sos24]. Yadrenko [Coh24]. Yihui [Joh20]. Ying [Par23]. York [Boo21, Fle22, Knu21, LL20, Par20].

Zero [HCS24]. Zero-Inflated [HCS24]. Zhang [Kim23, VW21, Zha21c].

References

Aikens:2023:ACP

[AB23] Rachael C. Aikens and Michael Baiocchi. Assignment-control plots: a visual companion for causal inference study design. The American Statistician, 77(1):72–84, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

            tronic). URL http://www.tandfonline.com/doi/full/10.
            1080/00031305.2022.2051605.

Ashner:2024:UIC

[AG24] Marissa C. Ashner and Tanya P. Garcia. Understanding the implications of a complete case analysis for regression models with a right-censored covariate. The American Statistician, 78 (3):335–344, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2282629.

Arnsbarger:2021:MRT

[AGK^{+}21] Madison Arnsbarger, Joshua Goldstein, Claire Kelling, Gizem Korkmaz, and Sallie Keller. Modeling response time to structure fires. The American Statistician, 75(1):92–100, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1695664.

Aldous:2021:PTP

[Ald21] David J. Aldous. A prediction tournament paradox. The American Statistician, 75(3):243–248, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2019.
1604430.

Alon:2023:HPD

[AM23] Noga Alon and Yaakov Malinovsky. Hitting a prime in 2.43 dice rolls (on average). The American Statistician, 77(3):301– 303, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2023.2179664.

Allen-Moyer:2022:IMV

[AMS22] Katherine Allen-Moyer and Jonathan Stallrich. Incorporating minimum variances into weighted optimality criteria. The American Statistician, 76(3):262–269, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2021.
1947375.

Andersson:2023:WCI

[And23] Per Gösta Andersson. The Wald confidence interval for a binomial p as an illuminating “Bad” example. The American Statistician, 77(4):443–448, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2183257.

Anonymous:2020:C

[Ano20a] Anonymous. Correction. The American Statistician, 74(2):212, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1708461.

Anonymous:2020:EC

[Ano20b] Anonymous. Editorial collaborators. The American Statistician, 74(4):420–421, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1842019.

Anonymous:2022:IAJ

[Ano22] Anonymous. The impact of application of the jackknife to the sample median. The American Statistician, 76(2):201, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2032827.

Anonymous:2023:CLU

[Ano23] Anonymous. Correction: Linearity of unbiased linear model estimators. The American Statistician, 77(2):237, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2184423. See [Por22].

Anonymous:2024:ASA

[Ano24] Anonymous. The American Statistician 2023 Associate Editors. The American Statistician, 78(1):??, 2024. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/00031305.
2024.2304534.

Arnold:2020:FER

[Arn20] Barry C. Arnold. Further examples related to the identical distribution of X/(X +Y) and Y/(X +Y). The American Statistician, 74(1):93–97, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2019.1575772.

Allen:2024:HMM

[AW24] Jessica Allen and Ting Wang. Hidden Markov models for lowfrequency earthquake recurrence. The American Statistician, 78 (1):100–110, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2282631.

Barker:2022:BRD

[Bar22] Chris Barker. Book review: Data Monitoring Committees in Clinical Trials: a Practical Perspective, 2nd ed. Susan S. Ellenberg, Thomas R. Fleming, and David L. DeMets, Hoboken, NJ, Wiley, 2019, xxii + 468 pp., $122.95(H), ISBN 978-1-119-51265-3. The American Statistician, 76(3):305–306, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2088199.

Benjamin:2019:TRI

[BB19] Daniel J. Benjamin and James O. Berger. Three recommendations for improving the use of p-values. The American Statistician, 73(S1):186–191, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1543135.

See [Fou20].

Berk:2021:ALR

[BBB^{+}21] Richard Berk, Andreas Buja, Lawrence Brown, Edward George, Arun Kumar Kuchibhotla, Weijie Su, and Linda Zhao. Assumption lean regression. The American Statistician, 75(1):76–84, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1592781.

Bower:2023:RRL

[BC23] Roy Bower and William Cipolli III. A response to Rice and Lumley. The American Statistician, 77(2):221–222, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2182362.

Boos:2021:PCU

[BD21] Dennis D. Boos and Siyu Duan. Pairwise comparisons using ranks in the one-way model. The American Statistician, 75(4):

414–423, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1860819.

Besse:2022:SBM

[BdBG^{+}22] Philippe Besse, Eustasio del Barrio, Paula Gordaliza, Jean-Michel Loubes, and Laurent Risser. A survey of bias in machine learning through the prism of statistical parity. The American Statistician, 76(2):188–198, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1952897.

Benoit:2021:BRT

[Ben21] Kenneth R. Benoit. Book review: Textual Data Science with R, Mónica Bécue-Bertaut. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xvii + 194 pp., $79.95 (H), $54.95 (e-book), ISBN: 978-1-138-62691-1 (H), 978-1-315-21266-1 (ebook). The American Statistician, 75(4):453–454, 2021. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.1985864.

Berger:2021:MHD

[Ber21] Roger L. Berger. McCann and habiger (2020), “the detection of nonnegligible directional effects with associated measures of statistical significance,”. The American Statistician, 75(1):115, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1850523.

Banks:2021:SCA

[BH21] David L. Banks and Mevin B. Hooten. Statistical challenges in agent-based modeling. The American Statistician, 75(3): 235–242, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.1900914.

Bower:2023:CN

[BHC^{+}23] Roy Bower, Justin Hager, Chris Cherniakov, Samay Gupta, and William Cipolli III. A case for nonparametrics. The American Statistician, 77(2):212–219, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2141858.

Bickel:2021:NHSb

[Bic21a] David R. Bickel. Null hypothesis significance testing defended and calibrated by Bayesian model checking. The American Statistician, 75(3):249–255, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1699443.

Bickel:2021:NHSa

[Bic21b] David R. Bickel. Null hypothesis significance testing interpreted and calibrated by estimating probabilities of sign errors: A Bayes-frequentist continuum. The American Statistician, 75(1): 104–112, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1816214.

Bilder:2022:ASC

[Bil22] Christopher R. Bilder. Alpha seminar: a course for new graduate students in statistics. The American Statistician, 76(3): 286–291, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2049366.

Bracher:2024:LFP

[BKKL24] Johannes Bracher, Nils Koster, Fabian Krüger, and Sebastian Lerch. Learning to forecast: The probabilistic time series forecasting challenge. The American Statistician, 78(1):115–127, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2023.2199800.

Bottai:2022:OCB

[BKL^{+}22] Matteo Bottai, Taeho Kim, Benjamin Lieberman, George Luta, and Edsel Peña. On optimal correlation-based prediction. The American Statistician, 76(4):313–321, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.

2051604. See comment [Lip23].

Blackburn:2020:BSS

[Bla20] McKinley L. Blackburn. Bias in small-sample inference with count-data models. The American Statistician, 74(3):267–273, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2018.1564699.

Blanc:2024:P

[Bla24] Adriana Verónica Blanc. The phistogram. The American Statistician, 78(2):229–239, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2267639.

Bartroff:2023:OFC

[BLW23a] Jay Bartroff, Gary Lorden, and Lijia Wang. Optimal and fast confidence intervals for hypergeometric successes. The American Statistician, 77(2):151–159, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.

2128421. See comment [Sch23] and response [BLW23b].

Bartroff:2023:RCS

[BLW23b] Jay Bartroff, Gary Lorden, and Lijia Wang. Response to comment by Schilling. The American Statistician, 77(3):343–344, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2023.2205455.

Bhagwat:2020:PBI

[BM20] Pankaj Bhagwat and Éric Marchand. On a proper Bayes, but inadmissible estimator. The American Statistician, 74(3):294– 296, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2019.1604432.

Boing-Messing:2021:BFT

[BMM21] Florian Böing-Messing and Joris Mulder. Bayes factors for testing order constraints on variances of dependent outcomes. The American Statistician, 75(2):152–161, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1715257.

Brown:2021:SSF

[BMS21] D. Andrew Brown, Christopher S. McMahan, and Stella Watson Self. Sampling strategies for fast updating of Gaussian Markov random fields. The American Statistician, 75(1):52–65, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1595144.

Boone:2021:BRM

[Boo21] Edward L. Boone. Book review: The Model Thinker: What You Need to Know to Make Data Work for You, by Scott E. Page. New York: Basic Books, 2018, xiii + 427 pp., $32.00 (hardback), ISBN: 978-0-465-09462-2. The American Statistician, 75(2):230–231, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1907993.

Baek:2023:MNC

[BP23] Jangsun Baek and Jeong-Soo Park. Mixture of networks for clustering categorical data: a penalized composite likelihood approach. The American Statistician, 77(3):259–273, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2141856.

Baldi:2020:BC

[BS20a] Pierre Baldi and Babak Shahbaba. Bayesian causality. The American Statistician, 74(3):249–257, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2019.
1647876.

Butcher:2020:BRF

[BS20b] Brandon Butcher and Brian J. Smith. Book review: Feature Engineering and Selection: A Practical Approach for Predictive Models, by Max Kuhn and Kjell Johnson. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xv + 297 pp., $79.95(H), ISBN: 978-1-138-07922-9. The American Statistician, 74(3):308–309, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1790217.

Bussberg:2021:STS

[Bus21] Nicholas W. Bussberg. Spatio-temporal statistics with r. The American Statistician, 75(1):114, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1865066.

Brini:2024:MDI

[BvdH24] Alberto Brini and Edwin R. van den Heuvel. Missing data imputation with high-dimensional data. The American Statistician, 78(2):240–252, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2259962.

Biro:2024:PCS

[BW24] Preston Biro and Stephen G. Walker. Play call strategies and modeling for target outcomes in football. The American Statistician, 78(1):66–75, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2223582.

Binette:2023:EPE

[BYH^{+}23] Olivier Binette, Sokhna A. York, Emma Hickerson, Youngsoo Baek, Sarvo Madhavan, and Christina Jones. Estimating the performance of entity resolution algorithms: Lessons learned through PatentsView.org. The American Statistician, 77(4): 370–380, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2191664.

Cheng:2024:ISM

[CC24] Nianpin Cheng and Beth Chance. Book review: Introduction to Statistical Modelling and Inference, Murray Aitkin, Boca Raton, FL: Chapman & Hall/CRC Press, 2023, xvi + 374 pp., $110.00(H), ISBN 978-1-032-10571-0. The American Statistician, 78(3):376–377, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2024.2326662.

Chen:2024:MMB

[CCL^{+}24] Chixiang Chen, Shuo Chen, Qi Long, Sudeshna Das, and Ming Wang. Multiple-model-based robust estimation of causal treatment effect on a binary outcome with integrated information from secondary outcomes. The American Statistician, 78(2): 150–160, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2250399.

Chattopadhyay:2024:OSW

[CCZ24] Ambarish Chattopadhyay, Eric R. Cohn, and José R. Zubizarreta. One-step weighting to generalize and transport treatment effect estimates to a target population. The American Statistician, 78(3):280–289, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2267598.

Cui:2022:BRS

[CF22] Jie Cui and Haoda Fu. Book review: Statistical Issues in Drug Development, 3rd ed., Stephen Senn, Hoboken, NJ: John Wiley and Sons, Ltd., 2021, xxi + 616 pp., $120.00 (H), ISBN: 978-1-119-23857-7. The American Statistician, 76(4):431, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2126685.

Campbell:2023:BFP

[CG23] Harlan Campbell and Paul Gustafson. Bayes factors and posterior estimation: Two sides of the very same coin. The American Statistician, 77(3):248–258, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2139293.

Charvat:2023:ULF

[Cha23] Hadrien Charvat. Using the Lambert function to estimate shared frailty models with a normally distributed random intercept. The American Statistician, 77(1):41–50, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2110939.

Chen:2020:MIP

[Che20] Qixuan Chen. Multiple imputation in practice: With examples using IVEware. The American Statistician, 74(4):417, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1831809.

Chen:2021:BRA

[Che21] Yen-Chi Chen. Book review: Applied Directional Statistics: Modern Methods and Case Studies, Christophe Ley and Thomas Verdebout, Boca Raton, FL, Chapman & Hall/CRC Press, 2018, xvi + 299 pp., $140.00 (hardback), ISBN: 978-1-138-62643-0. The American Statistician, 75(3):354, 2021. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.1949931.

Chen:2023:BRE

[Che23] Ding-Geng Chen. Book review: Event History Analysis with R, 2nd ed. The American Statistician, 77(3):340–341, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2230758.

Chern:2024:HPR

[Che24] Shane Chern. Hitting a prime by rolling a die with infinitely many faces. The American Statistician, 78(3):297–303, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2023.2290720.

Christensen:2020:CTT

[Chr20] Ronald Christensen. Comment on “Test for Trend With a Multinomial Outcome” by Szabo (2019). The American Statistician, 74(3):313–314, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1763835. See [Sza19].

Christensen:2022:COC

[Chr22a] Ronald Christensen. Comment on “On Optimal Correlation-Based Prediction,” by Bottai et al. (2022). The American Statistician, 76(4):322, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2055644. See [BKL^{+}22].

Christensen:2022:CPT

[Chr22b] Ronald Christensen. Comment on “On the Power of the F -test for Hypotheses in a Linear Model,” by Griffiths and Hill (2022). The American Statistician, 76(3):310–311, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2074541. See [GH22a].

Christensen:2024:CFK

[Chr24] Ronald Christensen. Comment on “Forbidden Knowledge and Specialized Training: A Versatile Solution for the Two Main Sources of Overfitting in Linear Regression,” by Rohlfs (2023). The American Statistician, 78(1):131–133, 2024. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2277156. See [Roh23].

Cinelli:2023:BRF

[Cin23] Carlos Cinelli. Book review: A First Course in Linear Model Theory, 2nd ed. The American Statistician, 77(4):451, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2261819.

Chatterjee:2020:SNA

[CMB20] Debashis Chatterjee, Trisha Maitra, and Sourabh Bhattacharya. A short note on almost sure convergence of Bayes factors in the general set-up. The American Statistician, 74(1):17–20, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1397548.

Cohen:2024:FPT

[Coh24] Joel E. Cohen. First-passage times for random partial sums: Yadrenko’s model for e and beyond. The American Statistician, 78(1):111–114, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2244542.

Carella:2022:BSA

[CTÁM22] Giulia Carella, Javier Pérez Trufero, Miguel Álvarez, and Jorge Mateu. A Bayesian spatial analysis of the heterogeneity in human mobility changes during the first wave of the COVID-19 epidemic in the United States. The American Statistician, 76 (1):64–72, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.1965657.

Christensen:2022:WYP

[CZ22] William F. Christensen and Brinley N. Zabriskie. When your permutation test is doomed to fail. The American Statistician, 76(1):53–63, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1902856.

DeOliveira:2020:MGB

[De 20] Victor De Oliveira. Models for geostatistical binary data: Properties and connections. The American Statistician, 74(1):72– 79, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2018.1444674.

Dai:2022:SM

[DH22a] Weixiao Dai and Toshimitsu Hamasaki. Statistics in medicine. The American Statistician, 76(2):199–200, 2022. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/00031305.
2022.2054626.

Dogucu:2022:CSU

[DH22b] Mine Dogucu and Jingchen Hu. The current state of undergraduate Bayesian education and recommendations for the future. The American Statistician, 76(4):405–413, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2089232.

deMast:2023:APS

[dMSNK23] Jeroen de Mast, Stefan H. Steiner, Wim P. M. Nuijten, and Daniel Kapitan. Analytical problem solving based on causal, correlational and deductive models. The American Statistician, 77 (1):51–61, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.2023633.

Davenport:2020:RPI

[DND^{+}20] Ernest C. Davenport, Jr., Kyle Nickodem, Mark L. Davison, Gareth Phillips, and Edmund Graham. The relative performance index: Neutralizing Simpson’s Paradox. The American Statistician, 74(2):116–124, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2018.1451777.

Das:2020:NEC

[DP20] Srinjoy Das and Dimitris N. Politis. Nonparametric estimation of the conditional distribution at regression boundary points. The American Statistician, 74(3):233–242, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1558109.

dosSantos:2023:SCW

[dSDUOG23] Marcelo dos Santos, Fernanda De Bastiani, Miguel A. Uribe-Opazo, and Manuel Galea. Selection criterion of working correlation structure for spatially correlated data. The American Statistician, 77(3):283–291, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2157874.

Epifanio:2020:AAM

[EIS20] Irene Epifanio, M. Victoria Ibáñez, and Amelia Simó. Archetypal analysis with missing data: See all samples by looking at a few based on extreme profiles. The American Statistician, 74 (2):169–183, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2018.1545700.

Elmore:2022:BCS

[EM22] Ryan Elmore and Gregory J. Matthews. Bang the can slowly: An investigation into the 2017 Houston Astros. The American Statistician, 76(2):110–116, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1902391.

Fosdick:2020:BRM

[FA20] Bailey K. Fosdick and G. Brooke Anderson. Book review: Modern Statistics for Modern Biology, by Susan Holmes and Wolfgang Huber. Cambridge, UK: Cambridge University Press, 2019, xxiii + 382 pp., $64.99(P), ISBN: 978-1-108-70529-5. The American Statistician, 74(3):309–311, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1790218.

Franck:2020:ABF

[FG20] Christopher T. Franck and Robert B. Gramacy. Assessing Bayes factor surfaces using interactive visualization and computer surrogate modeling. The American Statistician, 74(4): 359–369, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2019.1671219.

Fairley:2021:BES

[FH21] William B. Fairley and William A. Huber. On being an ethical statistical expert in a legal case. The American Statistician, 75(3):323–333, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1763834.

Far:2023:CSD

[FIP^{+}23] Serveh Sharifi Far, Vanda Inácio, Daniel Paulin, Miguel de Carvalho, Nicole H. Augustin, Mike Allerhand, and Gail Robertson. Consultancy style dissertations in statistics and data science: Why and how. The American Statistician, 77(3):331–339, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2163689.

Fisher:2021:VCT

[FK21] Aaron Fisher and Edward H. Kennedy. Visually communicating and teaching intuition for influence functions. The American Statistician, 75(2):162–172, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1717620.

Flegal:2021:DVC

[Fle21a] James M. Flegal. Data visualization: Charts, maps, and interactive graphics. Robert Grant. The American Statistician, 75(1):113, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1865062.

Flegal:2021:FPS

[Fle21b] James M. Flegal. Fundamentals of probability with stochastic processes, 4th ed. saeed ghahramani. The American Statistician, 75(1):113–114, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1865063.

Flegal:2022:BRD

[Fle22] James M. Flegal. Book review: Do Dice Play God? The Mathematics of Uncertainty, by Ian Stewart, New York, NY: Basic Books, 2019, 292 pp., $28.00 (H), ISBN: 978-1-5416-9947-2. The American Statistician, 76(1):85, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.2019999.

Flandre:2021:STL

[FO21] Philippe Flandre and John O’Quigley. The short-term and longterm hazard ratio model: Parameterization inconsistency. The American Statistician, 75(4):376–382, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1740786.

Foulley:2020:BDJ

[Fou20] Jean-Louis Foulley. Benjamin, D. J., and Berger, J. O. (2019), “Three Recommendations for Improving the Use of p-Values”, The American Statistician, 73, 186–191: Comment by Foulley. The American Statistician, 74(1):101–102, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1668850. See [BB19].

Freeman:2021:FAP

[Fre21] Peter E. Freeman. Facilitating authentic practice for early undergraduate statistics students. The American Statistician, 75 (4):433–444, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1844293.

Friendly:2020:VTE

[FS20] Michael Friendly and Matthew Sigal. Visualizing tests for equality of covariance matrices. The American Statistician, 74(2): 144–155, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2018.1497537.

Fecker:2021:CSB

[FSOW21] Julian Fecker, Martin Schumacher, Kristin Ohneberg, and Martin Wolkewitz. Correction of survival bias in a study about increased mortality of heads of government. The American Statistician, 75(1):85–91, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1638831.

Frey:2024:MCI

[FZ24] Jesse Frey and Yimin Zhang. Melded confidence intervals do not provide guaranteed coverage. The American Statistician, 78 (3):273–279, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2257253.

Gallagher:2020:WQA

[Gal20] Eugene D. Gallagher. Was Quetelet’s average man normal? The American Statistician, 74(3):301–306, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1706635.

Grimshaw:2020:GVB

[GBB20] Scott D. Grimshaw, Natalie J. Blades, and Candace Berrett. Going viral, binge-watching, and attention cannibalism. The American Statistician, 74(4):380–391, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1774415.

Graffelman:2023:IAV

[GdL23] Jan Graffelman and Jan de Leeuw. Improved approximation and visualization of the correlation matrix. The American Statistician, 77(4):432–442, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2186952.

Gelman:2023:TTL

[Gel23] Andrew Gelman. “Two truths and a lie” as a class-participation activity. The American Statistician, 77(1):97–101, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2058612.

Greenland:2020:CIP

[GFB^{+}20] Sander Greenland, Michael P. Fay, Erica H. Brittain, Joanna H. Shih, Dean A. Follmann, Erin E. Gabriel, and James M. Robins. On causal inferences for personalized medicine: How hidden causal assumptions led to erroneous causal claims about the D-value. The American Statistician, 74(3):243–248, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1575771.

Grantham:2020:TPC

[GFHK20] Kelsey L. Grantham, Andrew B. Forbes, Stephane Heritier, and Jessica Kasza. Time parameterizations in cluster randomized trial planning. The American Statistician, 74(2):184–189, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1623072.

Gu:2022:RKM

[GFY22] Jiaqi Gu, Yiwei Fan, and Guosheng Yin. Reconstructing the Kaplan–Meier estimator as an M-estimator. The American Statistician, 76(1):37–43, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2021.
1947376.

Ghosh:2020:LRL

[GGK20] Tamal Ghosh, Malay Ghosh, and Tatsuya Kubokawa. On the loss robustness of least-square estimators. The American Statistician, 74(1):64–67, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1529626.

Gomez:2020:CNC

[GGP20] Román Salmerón Gómez, Catalina García García, and Jose García Pérez. Comment on “A Note on Collinearity Diagnostics and Centering” by Velilla (2018). The American Statistician, 74(1): 68–71, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2019.1635527.

Griffiths:2022:PTH

[GH22a] William E. Griffiths and R. Carter Hill. On the power of the F -test for hypotheses in a linear model. The American Statistician, 76(1):78–84, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1979652. See [Har22b,

Chr22b] and rejoinder [GH22b].

Griffiths:2022:RHC

[GH22b] William E. Griffiths and R. Carter Hill. Rejoinder to Harville (2022) and Christensen (2022). Comments on “On the Power of the F -test for Hypotheses in a Linear Model,” by Griffiths and Hill (2022). The American Statistician, 76(3):312, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2074542. See [GH22a, Har22b, Chr22b].

Guo:2023:DPP

[GH23] Shijie Guo and Jingchen Hu. Data privacy protection and utility preservation through Bayesian data synthesis: a case study on Airbnb listings. The American Statistician, 77(2):192–200, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2077440.

Gelman:2024:CQD

[GHK24] Andrew Gelman, Jessica Hullman, and Lauren Kennedy. Causal quartets: Different ways to attain the same average treatment effect. The American Statistician, 78(3):267–272, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2267597.

Ghosh:2020:RJE

[Gho20] Malay Ghosh. Revisiting Jeffreys’ example: Bayes test of the normal mean. The American Statistician, 74(4):413–415, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1687013.

Gronau:2020:IBT

[GLW20] Quentin F. Gronau, Alexander Ly, and Eric-Jan Wagenmakers. Informed Bayesian t-tests. The American Statistician, 74 (2):137–143, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2018.1562983.

Gonzalez-Ortega:2021:HTP

[GOIRS21] J. González-Ortega, D. Ríos Insua, F. Ruggeri, and R. Soyer. Hypothesis testing in presence of adversaries. The American Statistician, 75(1):31–40, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1630001.

Gramacy:2020:SUO

[Gra20] Robert B. Gramacy. A shiny update to an old experiment game. The American Statistician, 74(1):87–92, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1505659.

Grabchak:2023:HDW

[Gra23] Michael Grabchak. How do we perform a paired t-test when we don’t know how to pair? The American Statistician, 77 (2):127–133, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2115552.

Goeman:2022:CTG

[GS22] Jelle J. Goeman and Aldo Solari. Comparing three groups. The American Statistician, 76(2):168–176, 2022. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/00031305.
2021.2002188.

Gajewski:2022:LEC

[GWJ22] Byron J. Gajewski, Jo A. Wick, and Truman J. Milling Jr. Letter to the Editor: A connection between baseball and clinical trials found in “Slugging percentage is not a percentage — and why that matters”. The American Statistician, 76(1):89, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.1990128.

Ge:2024:EIF

[GZWL24] Lin Ge, Yuzi Zhang, Lance A. Waller, and Robert H. Lyles. Enhanced inference for finite population sampling-based prevalence estimation with misclassification errors. The American Statistician, 78(2):192–198, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2250401.

Gu:2023:BLR

[GZY23] Jiaqi Gu, Yan Zhang, and Guosheng Yin. Bayesian logrank test. The American Statistician, 77(3):292–300, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2161637.

Hahn:2023:BRB

[Hah23] P. Richard Hahn. Book review: Bayesian Modeling and Computation in Python. The American Statistician, 77(4):450–451, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2023.2261818.

Hanley:2020:LWF

[Han20] James A. Hanley. Lest we forget: U.S. Selective Service Lotteries, 1917–2019. The American Statistician, 74(2):197–206, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1699444.

Harville:2022:BIU

[Har22a] David A. Harville. Bayesian inference is unaffected by selection: Fact or fiction? The American Statistician, 76(1):22–28, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1858963.

Harville:2022:CPT

[Har22b] David A. Harville. Comment on “On the Power of the F - test for Hypotheses in a Linear Model,” by Griffiths and Hill (2022). The American Statistician, 76(3):308–309, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2074540. See [GH22a].

Han:2024:SAC

[HAT24] Larry Han, Andrea Arfè, and Lorenzo Trippa. Sensitivity analyses of clinical trial designs: Selecting scenarios and summarizing operating characteristics. The American Statistician, 78 (1):76–87, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2216253.

Hanley:2022:PDH

[HB22a] James A. Hanley and Sahir Bhatnagar. The “Poisson” distribution: History, reenactments, adaptations. The American Statistician, 76(4):363–371, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2046159.

Hui:2022:SCG

[HB22b] Francis K. C. Hui and Howard D. Bondell. Spatial confounding in generalized estimating equations. The American Statistician, 76(3):238–247, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.2009372.

Hwang:2024:CUE

[HCS24] Wen-Han Hwang, Lu-Fang Chen, and Jakub Stoklosa. Counting the unseen: Estimation of susceptibility proportions in zeroinflated models using a conditional likelihood approach. The American Statistician, 78(2):161–170, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2023.
2249529.

Hines:2022:DSL

[HDDOV22] Oliver Hines, Oliver Dukes, Karla Diaz-Ordaz, and Stijn Vansteelandt. Demystifying statistical learning based on efficient influence functions. The American Statistician, 76(3):292–304, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.2021984.

Hering:2021:IRS

[HDM21] Amanda S. Hering, Luke Durell, and Grant Morgan. Illustrating randomness in statistics courses with spatial experiments. The American Statistician, 75(3):343–353, 2021. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1871070.

Hemerik:2024:TRT

[Hem24] Jesse Hemerik. On the term “Randomization test”. The American Statistician, 78(3):327–334, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2024.
2319182.

Hackenberg:2022:UDP

[HGK^{+}22] Maren Hackenberg, Marlon Grodd, Clemens Kreutz, Martina Fischer, Janina Esins, Linus Grabenhenrich, Christian Karagiannidis, and Harald Binder. Using differentiable programming for flexible statistical modeling. The American Statistician, 76(3): 270–279, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.2002189.

Higgins:2021:OEM

[HHL21] James J. Higgins, Michael J. Higgins, and Jinguang Lin. From one environment to many: The problem of replicability of statistical inferences. The American Statistician, 75(3):334–342, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1829047.

Hitchcock:2024:LDB

[Hit24] David B. Hitchcock. Lessons from a discussion-based course on the history of statistics. The American Statistician, 78(3): 368–374, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2281359.

Hooten:2021:MRB

[HJB21] Mevin B. Hooten, Devin S. Johnson, and Brian M. Brost. Making recursive Bayesian inference accessible. The American Statistician, 75(2):185–194, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1665584.

Hu:2020:IEC

[HJQ20a] Xinjie Hu, Aekyung Jung, and Gengsheng Qin. Interval estimation for the correlation coefficient. The American Statistician, 74(1):29–36, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.
com/doi/full/10.1080/00031305.2018.1437077. See comment

[KX20] and response [HJQ20b].

Hu:2020:RLE

[HJQ20b] Xinjie Hu, Aekyung Jung, and Gengsheng Qin. A response to the letter to the Editor on “Interval Estimation for the Correlation Coefficient,” The American Statistician, 74:1, 29-36: Comment by Krishnamoorthy and Xia. The American Statistician, 74(4): 419, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2020.1827032. See [HJQ20a, KX20].

Habiger:2022:PPR

[HL22] Joshua Habiger and Ye Liang. Publication policies for replicable research and the community-wide false discovery rate. The American Statistician, 76(2):131–141, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2021.
1999857.

Hoaglin:2020:LED

[Hoa20] David C. Hoaglin. Letter to the Editor: Did Phlegon actually use a stem-and-leaf display? The American Statistician, 74(2):211, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1721329.

Hofert:2021:RNG

[Hof21] Marius Hofert. Random number generators produce collisions: Why, how many and more. The American Statistician, 75(4): 394–402, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1782261.

Hofert:2023:RPP

[HPZ23] Marius Hofert, Avinash Prasad, and Mu Zhu. RafterNet: Probabilistic predictions in multi-response regression. The American Statistician, 77(4):406–416, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2141857.

Hansen:2023:CTI

[HR23] Spencer Hansen and Ken Rice. Coherent tests for interval null hypotheses. The American Statistician, 77(1):20–28, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2050299.

Hardwicke:2023:SGA

[HSHM^{+}23] Tom E. Hardwicke, Maia Salholz-Hillel, Mario Malicki, Dénes Szücs, Theiss Bendixen, and John P. A. Ioannidis. Statistical guidance to authors at top-ranked journals across scientific disciplines. The American Statistician, 77(3):239–247, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2143897.

Han:2023:BFH

[HSLY23] Gang Han, Thomas J. Santner, Haiqun Lin, and Ao Yuan. Bayesian-frequentist hybrid inference in applications with small sample sizes. The American Statistician, 77(2):143–150, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2127897.

Huang:2023:AUD

[Hua23] Alan Huang. On arbitrarily underdispersed discrete distributions. The American Statistician, 77(1):29–34, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2106305.

Hawinkel:2024:SEI

[HWM24] Stijn Hawinkel, Willem Waegeman, and Steven Maere. Out-of-sample R^{2}: Estimation and inference. The American Statistician, 78(1):15–25, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2216252.

Hutson:2023:STP

[HY23] Alan D. Hutson and Han Yu. The sign test, paired data, and asymmetric dependence: a cautionary tale. The American Statistician, 77(1):35–40, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2110938.

Han:2023:SPP

[HYW23] Zifei Han, Keying Ye, and Min Wang. A study on the power parameter in power prior Bayesian analysis. The American Statistician, 77(1):12–19, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2046160.

Johnson:2020:HDS

[JDC20] Tim Johnson, Christopher T. Dawes, and Dalton Conley. How does a statistician raise an army? The time when John W. Tukey, a team of luminaries, and a statistics graduate student repaired the Vietnam Selective Service Lotteries. The American Statistician, 74(2):190–196, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1677267.

Johansson:2022:IEC

[JN22] Per Johansson and Mattias Nordin. Inference in experiments conditional on observed imbalances in covariates. The American Statistician, 76(4):394–404, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2054859.

Johnstone:2024:UCW

[JN24] Chancellor Johnstone and Dan Nettleton. Using conformal win probability to predict the winners of the canceled 2020 NCAA basketball tournaments. The American Statistician, 78(3):304– 317, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2023.2283199.

Joshi:2024:PBD

[JNCP24] Chaitanya Joshi, Charné Nel, Javier Cano, and Devon L. L. Polaschek. Parole board decision-making using adversarial risk analysis. The American Statistician, 78(3):345–358, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2024.2303416.

Johnson:2020:BRR

[Joh20] Paul Johnson. Book review: R Markdown: The Definitive Guide, Yihui Xie, J. J. Allaire, and Garrett Grolemund. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xxxiv + 303 pp., $38.95(P), ISBN: 978-1-138-35933-8. The American Statistician, 74(2):209–210, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1745577.

Jiang:2021:SEM

[JSHZ21] Wei Jiang, Shuang Song, Lin Hou, and Hongyu Zhao. A set of efficient methods to generate high-dimensional binary data with specified correlation structures. The American Statistician, 75(3):310–322, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1816213.

Kaplan:2023:IEL

[KB23] Andee Kaplan and Jacob Bien. Interactive exploration of large dendrograms with prototypes. The American Statistician, 77 (2):201–211, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2087734.

Kranker:2021:IEE

[KBF21] Keith Kranker, Laura Blue, and Lauren Vollmer Forrow. Improving effect estimates by limiting the variability in inverse propensity score weights. The American Statistician, 75(3): 276–287, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1737229.

Kaplan:2022:PAP

[KBS22] Andee Kaplan, Brenda Betancourt, and Rebecca C. Steorts. A practical approach to proper inference with linked data. The American Statistician, 76(4):384–393, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2041482.

Kejzlar:2023:BBV

[KBSM23] Vojtech Kejzlar, Shrijita Bhattacharya, Mookyong Son, and Tapabrata Maiti. Black Box variational Bayesian model averaging. The American Statistician, 77(1):85–96, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2058611.

Kunzmann:2021:RBP

[KGL^{+}21] Kevin Kunzmann, Michael J. Grayling, Kim May Lee, David S. Robertson, Kaspar Rufibach, and James M. S. Wason. A review of Bayesian perspectives on sample size derivation for confirmatory trials. The American Statistician, 75(4):424–432, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.1901782.

Kejzlar:2024:IVI

[KH24] Vojtech Kejzlar and Jingchen Hu. Introducing variational inference in statistics and data science curriculum. The American Statistician, 78(3):359–367, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2023.
2232006.

Kim:2020:DTH

[Kim20a] Jae H. Kim. Decision-theoretic hypothesis testing: A primer with R package OptSig. The American Statistician, 74(4):370– 379, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2020.1750484.

Kim:2020:BRP

[Kim20b] Yongdai Kim. Book review: The 9 Pitfalls of Data Science, by Gary Smith and Jay Cordes. Oxford, United Kingdom: Oxford University Press, 2019, v + 256 pp., $32.95(H), ISBN: 978-0-19-884439-6. The American Statistician, 74(3):307, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1790216.

Kim:2023:BRG

[Kim23] Jae-Kwang Kim. Book review: Graph Sampling, Li-Chun Zhang, Boca Raton, FL: Chapman & Hall/CRC Press, 2022, xv + 121 pp., $59.95(H), ISBN: 978-1-032-06708-7. The American Statistician, 77(2):234, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2198354.

Kapelner:2021:HOD

[KKS^{+}21] Adam Kapelner, Abba M. Krieger, Michael Sklar, Uri Shalit, and David Azriel. Harmonizing optimized designs with classic randomization in experiments. The American Statistician, 75 (2):195–206, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1717619.

Kline:2022:BFB

[Kli22] Brendan Kline. Bayes factors based on p-values and sets of priors with restricted strength. The American Statistician, 76(3): 203–213, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.1877815.

Kane:2020:NAS

[KM20] Andrew Kane and Abhyuday Mandal. A new analysis strategy for designs with complex aliasing. The American Statistician, 74(3):274–281, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1585287.

Karanevich:2021:OSS

[KMG^{+}21] Alex Karanevich, Richard Meier, Stefan Graw, Anna Mc-Glothlin, and Byron Gajewski. Optimizing sample size allocation and power in a Bayesian two-stage drop-the-losers design. The American Statistician, 75(1):66–75, 2021. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1610065.

Knudson:2021:BRM

[Knu21] Christina P. Knudson. Book review: x + y: A Mathematician’s Manifesto for Rethinking Gender, by Eugenia Cheng. New York: Basic Books, 2020, x + 272 pp., $28.00 (hardback), ISBN: 978-1-5416-4650-6. The American Statistician, 75(2):232–233, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.1907998.

Kozubowski:2020:GMR

[KP20] Tomasz J. Kozubowski and Krzysztof Podgórski. Gaussian mixture representation of the Laplace distribution revisited: Bibliographical connections and extensions. The American Statistician, 74(4):407–412, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1630000.

Kross:2020:DDS

[KPC^{+}20] Sean Kross, Roger D. Peng, Brian S. Caffo, Ira Gooding, and Jeffrey T. Leek. The democratization of data science education. The American Statistician, 74(1):1–7, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1668849.

Keele:2021:CCP

[KS21] Luke Keele and Dylan S. Small. Comparing covariate prioritization via matching to machine learning methods for causal inference using five empirical applications. The American Statistician, 75(4):355–363, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1867638.

Kvam:2021:PRA

[Kva21] Paul Kvam. The price is right: Analyzing bidding behavior on contestants’ row. The American Statistician, 75(1):15–22, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1592782.

Krishnamoorthy:2020:LEX

[KX20] Kalimuthu Krishnamoorthy and Yanping Xia. Letters to the editor: Xinjie Hu, Aekyung Jung, and Gengsheng Qin (2020), “Interval Estimation for the Correlation Coefficient,” The American Statistician, 74:1, 29–36: Comment by Krishnamoorthy and Xia. The American Statistician, 74(4):418, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1829048. See [HJQ20a, HJQ20b].

Liu:2022:BAM

[LAMM22] Juxin Liu, Annshirley Afful, Holly Mansell, and Yanyuan Ma. Bias analysis for misclassification errors in both the response variable and covariate. The American Statistician, 76(4):353– 362, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2022.2066725.

Larsson:2024:CDA

[Lar24] Rolf Larsson. Confidence distributions for the autoregressive parameter. The American Statistician, 78(1):58–65, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2226184.

Liao:2021:RPB

[LBM21] J. G. Liao, Arthur Berg, and Timothy L. McMurry. A robustified posterior for Bayesian inference on a large number of parallel effects. The American Statistician, 75(2):145–151, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1701549.

Lee:2022:CPC

[LBV^{+}22] Shing Lee, Emilia Bagiella, Roger Vaughan, Usha Govindarajulu, Paul Christos, Denise Esserman, Hua Zhong, and Mimi Kim. COVID-19 pandemic as a change agent in the structure and practice of statistical consulting centers. The American Statistician, 76(2):152–158, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.2023045.

Li:2022:RAA

[LCHL22] Yao Li, Minhao Cheng, Cho-Jui Hsieh, and Thomas C. M. Lee. A review of adversarial attack and defense for classification methods. The American Statistician, 76(4):329–345, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.2006781.

Loux:2021:APE

[LD21] Travis Loux and Orlando Davy. Adjusting published estimates for exploratory biases using the truncated normal distribution. The American Statistician, 75(3):294–299, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1775700.

Lee:2021:BRB

[Lee21] Youjin Lee. Book review: Beyond Multiple Linear Regression: Applied Generalized Linear Models and Multilevel Models in R, by Paul Roback and Julie Legler. Boca Raton, FL: Chapman & Hall/CRC Press, 2020 (H)/2021 (e-book), xvii + 418 pp., $99.95 (H/e-book), ISBN: 978-1-4398-8538-3 (H), 978-0-429-06666-5 (e-book). The American Statistician, 75(4):450– 451, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2021.1985862.

Li:2023:CBM

[LGV23] Xun Li, Joyee Ghosh, and Gabriele Villarini. A comparison of Bayesian multivariate versus univariate normal regression models for prediction. The American Statistician, 77(3):304–312, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2087735.

Lavine:2021:IOC

[LH21] Michael Lavine and Jim Hodges. Intuition for an old curiosity and an implication for MCMC. The American Statistician, 75(1):1–6, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2018.1518267.

Li:2022:BAI

[Li22] Qiwei Li. Bayesian analysis of infectious diseases: COVID-19 and beyond. The American Statistician, 76(2):199, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2054625.

Ligtvoet:2021:EBF

[Lig21] Rudy Ligtvoet. Exact Bayes factors for the comparison of multinomial distributions. The American Statistician, 75(1):7–14, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2019.1575773.

Lipovetsky:2023:COC

[Lip23] Stan Lipovetsky. Comment on “On Optimal Correlation-Based Prediction”, By Bottai et al. (2022). The American Statistician, 77(1):113, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2141879. See [BKL^{+}22].

Li:2020:MSS

[LJ20] Wen Li and Thomas O. Jemielita. Mathematical and statistical skills in the biopharmaceutical industry: A pragmatic approach. The American Statistician, 74(4):416–417, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1831806.

Lubich:2022:SIM

[LJY22] Bradley Lubich, Daniel Jeske, and Weixin Yao. Statistical inference for method of moments estimators of a semi-supervised two-component mixture model. The American Statistician, 76 (4):376–383, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2096695.

Lee:2020:TRL

[LL20] Mintaek Lee and Jaechoul Lee. Trend and return level of extreme snow events in New York City. The American Statistician, 74(3):282–293, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1592780.

Long:2023:CCT

[LLZL23] Mingya Long, Zhengbang Li, Wei Zhang, and Qizhai Li. The Cauchy combination test under arbitrary dependence structures. The American Statistician, 77(2):134–142, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2116109.

Liao:2021:CCB

[LMB21] J. G. Liao, Vishal Midya, and Arthur Berg. Connecting and contrasting the Bayes factor and a modified ROPE procedure for testing interval null hypotheses. The American Statistician, 75 (3):256–264, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2019.1701550.

Lott:2020:WCI

[LR20] Anne Lott and Jerome P. Reiter. Wilson confidence intervals for binomial proportions with multiple imputation for missing data. The American Statistician, 74(2):109–115, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1473796.

Liu:2022:MII

[LR22] Bo Liu and Jerome P. Reiter. Multiple imputation inference with integer-valued point estimates. The American Statistician, 76 (4):323–328, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.2006780.

Larsen:2024:SCO

[LSS^{+}24] Nicholas Larsen, Jonathan Stallrich, Srijan Sengupta, Alex Deng, Ron Kohavi, and Nathaniel T. Stevens. Statistical challenges in online controlled experiments: A review of A/B testing methodology. The American Statistician, 78(2):135–149, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2023.2257237.

Lund:2020:BRT

[Lun20] Robert B. Lund. Book review: Time Series: A Data Analysis Approach Using R, Robert H. Shumway and David S. Stoffer. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xii + 259 pp., $79.95(H), ISBN: 978-0-367-22109-6. The American Statistician, 74(3):312, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1790221.

Leng:2020:CRC

[LZ20] Ling Leng and Wei Zhu. Compound regression and constrained regression: Nonparametric regression frameworks for EIV models. The American Statistician, 74(3):226–232, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1556734.

LaHaye:2021:LCC

[LZ21] Roberta La Haye and Petr Zizler. The Lorenz curve in the classroom. The American Statistician, 75(2):217–225, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1822916.

Liao:2024:PVO

[LZN^{+}24] Lauren D. Liao, Yeyi Zhu, Amanda L. Ngo, Rana F. Chehab, and Samuel D. Pimentel. Prioritizing variables for observational study design using the joint variable importance plot. The American Statistician, 78(3):318–326, 2024. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/00031305.
2024.2303419.

Mandel:2020:SUM

[Man20] Micha Mandel. The scaled uniform model revisited. The American Statistician, 74(1):98–100, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2019.1604431.

See comment [Tar20].

Marino:2024:BRA

[Mar24] Maria Francesca Marino. Book review: Applied Linear Regression for Longitudinal Data: With an Emphasis on Missing Observations, Frans E. S. Tan and Shahab Jolani, Boca Raton, FL: Chapman & Hall/CRC Press, 2023, xxi + 226 pp., $120.00(H), ISBN 978-0-367-63431-5. The American Statistician, 78(1):128–129, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2024.2302792.

MacDonald:2020:TSM

[MB20] Iain L. MacDonald and Feroz Bhamani. A time-series model for underdispersed or overdispersed counts. The American Statistician, 74(4):317–328, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2018.1505656.

McIntosh:2022:CSS

[McI22] Matthew J. McIntosh. Calculating sample size for Follmann’s simple multivariate test for one-sided alternatives. The American Statistician, 76(1):16–21, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1787224.

Moradi:2023:HST

[MCPGM23] Mehdi Moradi, Ottmar Cronie, Unai Pérez-Goya, and Jorge Mateu. Hierarchical spatio-temporal change-point detection. The American Statistician, 77(4):390–400, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2023.
2191670.

Matabuena:2023:HTM

[MFD^{+}23] Marcos Matabuena, Paulo Félix, Marc Ditzhaus, Juan Vidal, and Francisco Gude. Hypothesis testing for matched pairs with missing data by maximum mean discrepancy: An application to continuous glucose monitoring. The American Statistician, 77 (4):357–369, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2200512.

Miao:2024:ALR

[MG24] Weiwen Miao and Joseph L. Gastwirth. The application of the likelihood ratio test and the Cochran–Mantel–Haenszel test to discrimination cases. The American Statistician, 78(2):253–263, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2023.2259969.

Mansor:2020:DDT

[MGM20] Mahayaudin M. Mansor, David A. Green, and Andrew V. Metcalfe. Detecting directionality in time series. The American Statistician, 74(3):258–266, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2018.
1545699.

McCann:2020:DND

[MH20] Melinda H. McCann and Joshua D. Habiger. The detection of nonnegligible directional effects with associated measures of statistical significance. The American Statistician, 74(3):213–217, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2018.1497538. See response [MH21].

McCann:2021:RLE

[MH21] Melinda H. McCann and Joshua D. Habiger. Response to the letter to the Editor on “The Detection of Nonnegligible Directional Effects With Associated Measures of Statistical Significance,” The American Statistician, 74:3, 213–217: Comment by Roger Berger. The American Statistician, 75(1):116, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1851766. See [MH20].

Michael:2024:BRI

[Mic24] Skevi Michael. Book review: Introduction to Stochastic Finance with Market Examples, 2nd ed., Nicolas Privault, Boca Raton, FL: Chapman & Hall/CRC Press, 2023, x + 652 pp., $120.00(H), ISBN 978-1-032-28826-0. The American Statistician, 78(1):129–130, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2024.2303414.

Matabuena:2023:EKM

[MKR^{+}23] Marcos Matabuena, Marta Karas, Sherveen Riazati, Nick Caplan, and Philip R. Hayes. Estimating knee movement patterns of recreational runners across training sessions using multilevel functional regression models. The American Statistician, 77(2): 169–181, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2105950.

Motalebi:2021:HBS

[MSS21] Narges Motalebi, Nathaniel T. Stevens, and Stefan H. Steiner. Hurdle blockmodels for sparse network modeling. The American Statistician, 75(4):383–393, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1865199.

Mukhopadhyay:2022:PIM

[Muk22] Nitis Mukhopadhyay. Pairwise independence may not imply independence: New illustrations and a generalization. The American Statistician, 76(2):184–187, 2022. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2039763.

Mulder:2023:BTL

[Mul23] Joris Mulder. Bayesian testing of linear versus nonlinear effects using Gaussian process priors. The American Statistician, 77(1):1–11, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2028675.

Manrique-Vallier:2020:BRC

[MV20] Daniel Manrique-Vallier. Book review: Capture–Recapture Methods for the Social and Medical Sciences, Dankmar Böhning, Peter G. M. van der Heijden, and John Bunge, eds. Boca Raton, FL: Chapman & Hall/CRC Press, 2017, xxxii + 429 pp., $130.00(H), ISBN: 978-1-4987-4531-4. The American Statistician, 74(2):207–208, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1745574.

Mulder:2022:GSD

[MWM22] Joris Mulder, Eric-Jan Wagenmakers, and Maarten Marsman. A generalization of the Savage–Dickey density ratio for testing equality and order constrained hypotheses. The American Statistician, 76(2):102–109, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1799861.

McGee:2023:ARM

[MWS23] Monnie McGee, Benjamin Williams, and Jacy Sparks. Athlete recruitment and the myth of the sophomore peak. The American Statistician, 77(2):182–191, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2127896.

Nethery:2022:SIE

[NCK^{+}22] Rachel C. Nethery, Jarvis T. Chen, Nancy Krieger, Pamela D. Waterman, Emily Peterson, Lance A. Waller, and Brent A. Coull. Statistical implications of endogeneity induced by residential segregation in small-area modeling of health inequities. The American Statistician, 76(2):142–151, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.2003245.

Nguyen:2021:PAM

[Ngu21] Duy Nguyen. A probabilistic approach to the moments of binomial random variables and application. The American Statistician, 75(1):101–103, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1679257.

Ni:2022:BRE

[Ni22] Yang Ni. Book review: Exploratory Data Analysis with MAT-LAB, 3rd ed., by Wendy L. Martinez, Angel R. Martinez, and Jeffrey L. Solka. The American Statistician, 76(1):85–86, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.2020000.

Ni:2024:BRD

[Ni24] Yang Ni. Book review: Deep Learning and Scientific Computing with R torch, Sigrid Keydana, Boca Raton, FL: Chapman & Hall/CRC Press, 2023, xix + 393 pp., $180.00(H), ISBN: 978-1-032-23138-9. The American Statistician, 78(2):264, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2024.2320219.

Nassiri:2020:IMI

[NMVBB20] Vahid Nassiri, Geert Molenberghs, Geert Verbeke, and João Barbosa-Breda. Iterative multiple imputation: A framework to determine the number of imputed datasets. The American Statistician, 74(2):125–136, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2018.1543615.

Noguchi:2022:GDM

[NR22] Kimihiro Noguchi and Koby F. Robles. On generating distributions with the memoryless property. The American Statistician, 76(3):280–285, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.2006782.

Nguyen:2024:HCS

[NYM24] Quang Nguyen, Ronald Yurko, and Gregory J. Matthews. Here comes the STRAIN: Analyzing defensive pass rush in American football with player tracking data. The American Statistician, 78(2):199–208, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2242442.

Ogasawara:2020:SIM

[Oga20] Haruhiko Ogasawara. Some improvements on Markov’s theorem with extensions. The American Statistician, 74(3):218–225, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2018.1497539.

ONeill:2021:COD

[O’N21] Ben O’Neill. The classical occupancy distribution: Computation and approximation. The American Statistician, 75(4): 364–375, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2019.1699445.

Park:2020:BRA

[Par20] Jong Hee Park. Book review: The Art of Statistics: How to Learn From Data, David Spiegelhalter. New York, NY: Basic Books, 2019, xvi + 426 pp., $32.00(H), ISBN: 978-1-5416-1851-0. The American Statistician, 74(2):207, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1745572.

Park:2023:BRH

[Par23] Junyong Park. Book review: Handbook of Multiple Comparisons, Xinping Cui, Thorsten Dickhaus, Ying Ding, and Jason C. Hsu, eds., Boca Raton, FL: Chapman & Hall/ CRC Press, 2022, xiv + 404 pp., $250.00(H), ISBN: 978-0-367-14067-0. The American Statistician, 77(2):234–236, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2198355.

Petersen:2024:TVP

[PE24] Anne Helby Petersen and Claus Ekstrøm. Technical validation of plot designs by use of deep learning. The American Statistician, 78(2):220–228, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2270649.

Perkovic:2022:BRL

[Per22a] Emilija Perković. Book review: Leadership in Statistics and Data Science: Planning for Inclusive Excellence, by Amanda L. Golbeck, ed. Cham, Switzerland, Springer, 2021, xxvii + 432 pp., $159.99(H), ISBN: 978-3-030-60059-4. The American Statistician, 76(3):306–307, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2088201.

Perkovic:2022:BRP

[Per22b] Emilija Perković. Book review: The Phantom Pattern Problem: The Mirage of Big Data, Gary Smith and Jay Cordes. Oxford, United Kingdom: Oxford University Press, 2020, vii + 227 pp., $32.95 (H), ISBN: 978-0-19-886416-5. The American Statistician, 76(1):86–87, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.2020002.

Peskun:2020:TTV

[Pes20] Peter H. Peskun. Two-tailed p-values and coherent measures of evidence. The American Statistician, 74(1):80–86, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1475304.

Puig:2023:MLE

[PG23] Xavier Puig and Josep Ginebra. Mapping life expectancy loss in Barcelona in 2020. The American Statistician, 77(4):417–424, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2023.2197022.

Park:2022:SEP

[PGW22] Chanseok Park, Kun Gou, and Min Wang. A study on estimating the parameter of the truncated geometric distribution. The American Statistician, 76(3):257–261, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2034666.

Pandya:2024:BDB

[PLBM24] Sachin S. Pandya, Xiaomeng Li, Eric Barón, and Timothy E. Moore. Bayesian detection of bias in peremptory challenges using historical strike data. The American Statistician, 78(2): 209–219, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2249967.

Pawel:2024:ECC

[PLW24] Samuel Pawel, Alexander Ly, and Eric-Jan Wagenmakers. Evidential calibration of confidence intervals. The American Statistician, 78(1):47–57, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2216239.

Portnoy:2022:LUL

[Por22] Stephen Portnoy. Linearity of unbiased linear model estimators. The American Statistician, 76(4):372–375, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2076743. See [Ano23].

Presnell:2022:GDC

[Pre22] Brett Presnell. A geometric derivation of the Cantor distribution. The American Statistician, 76(1):73–77, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.1905062.

Papadopoulos:2021:DHH

[PS21] Alecos Papadopoulos and Roland B. Stark. Does home health care increase the probability of 30-day hospital readmissions? Interpreting coefficient sign reversals, or their absence, in binary logistic regression analysis. The American Statistician, 75(2): 173–184, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2019.1704873.

Paindaveine:2023:RNV

[PS23] Davy Paindaveine and Philippe Spindel. Revisiting the name variant of the two-children problem. The American Statistician, 77(4):401–405, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2173293.

Peng:2021:OBT

[PW21] Bo Peng and Min Wang. Objective Bayesian testing for the correlation coefficient under divergence-based priors. The American Statistician, 75(1):41–51, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2019.
1677266.

Qiu:2020:BDS

[Qiu20] Peihua Qiu. Big data? statistical process control can help! The American Statistician, 74(4):329–344, 2020. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/00031305.
2019.1700163.

Qin:2023:PMF

[QSL23] Yuxin Qin, Heather Sasinowska, and Lawrence Leemis. The probability mass function of the Kaplan–Meier product-limit estimator. The American Statistician, 77(1):102–110, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2070279.

Ramsay:2023:OOD

[Ram23] James O. Ramsay. Object oriented data analysis. The American Statistician, 77(1):111, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2160590.

Reiss:2021:PDJ

[Rei21] Philip T. Reiss. A problem of distributive justice, solved by the lasso. The American Statistician, 75(2):139–144, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1688682.

Rauwolf:2023:QIP

[RK23] Diana Rauwolf and Udo Kamps. Quantifying the inspection paradox with random time. The American Statistician, 77(3): 274–282, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2151510.

Rugamer:2024:SSD

[RKK24] David Rügamer, Chris Kolb, and Nadja Klein. Semi-structured distributional regression. The American Statistician, 78(1):88– 99, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2022.2164054.

Rice:2023:CCN

[RL23] Kenneth Rice and Thomas Lumley. Comment on “A Case for Nonparametrics” by Bower et al. The American Statistician, 77(2):220, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2172078.

Rohlfs:2023:FKS

[Roh23] Chris Rohlfs. Forbidden knowledge and specialized training: a versatile solution for the two main sources of overfitting in linear regression. The American Statistician, 77(2):160–168, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2128874. See comment [Chr24].

Rosenbaum:2022:NTT

[Ros22] Paul R. Rosenbaum. A new transformation of treated-control matched-pair differences for graphical display. The American Statistician, 76(4):346–352, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2063944.

Roths:2022:BRP

[Rot22] Scott A. Roths. Book review: Probability, Statistics, and Data: a Fresh Approach Using R, Darrin Speegle and Bryan Clair, Boca Raton, FL: Chapman & Hall/CRC Press, 2022, xii + 500 pp., $99.95(H), ISBN: 978-0-367-43667-4. The American Statistician, 76(4):430, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2126684.

Rougier:2021:EFO

[RP21] Jonathan Rougier and Carey E. Priebe. The exact form of the “Ockham Factor” in model selection. The American Statistician, 75(3):288–293, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1764865.

Raman:2023:IEG

[RUCH23] Rameela Raman, Jessica Utts, Andrew I. Cohen, and Matthew J. Hayat. Integrating ethics into the Guidelines for Assessment and Instruction in Statistics Education (GAISE). The American Statistician, 77(3):323–330, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2022.
2156612.

Rice:2022:ERU

[RY22] Kenneth Rice and Lingbo Ye. Expressing regret: a unified view of credible intervals. The American Statistician, 76(3): 248–256, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2039764.

Schilling:2023:BJL

[Sch23] Mark F. Schilling. Bartroff, J., Lorden, G. and Wang, L. (2022), “Optimal and Fast Confidence Intervals for Hypergeometric Successes,” The American Statistician: Comment by Schilling. The American Statistician, 77(3):342, 2023. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2197021. See [BLW23a] and response

[BLW23b].

Su:2020:NIR

[SCW20] Chien-Lang Su, Sun-Hao Chang, and Ruby Chiu-Hsing Weng. A note on item response theory modeling for online customer ratings. The American Statistician, 74(1):53–63, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1422804.

Sainsbury-Dale:2024:LFP

[SDZMH24] Matthew Sainsbury-Dale, Andrew Zammit-Mangion, and Raphaël Huser. Likelihood-free parameter estimation with neural Bayes estimators. The American Statistician, 78(1):1–14, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/00031305.
2023.2249522.

Segal:2021:TRC

[Seg21] Brian D. Segal. Toward replicability with confidence intervals for the exceedance probability. The American Statistician, 75 (2):128–138, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2019.1678521.

Sidi:2022:DBB

[SH22a] Yulia Sidi and Ofer Harel. Difference between binomial proportions using newcombe’s method with multiple imputation for incomplete data. The American Statistician, 76(1):29–36, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.1898468.

Stevens:2022:CPM

[SH22b] Nathaniel T. Stevens and Luke Hagar. Comparative probability metrics: Using posterior probabilities to account for practical equivalence in A/B tests. The American Statistician, 76 (3):224–237, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.2000495.

Shin:2020:BRM

[Shi20] Seung Jun Shin. Book review: Model-Based Clustering and Classification for Data Science: With Applications in R, Charles Bouveyron, Gilles Celeux, T. Brendan Murphy, and Adrian E. Raftery. Cambridge, UK: Cambridge University Press, 2019, xvii + 427 pp., $79.99(H), ISBN: 978-1-108-49420-5. The American Statistician, 74(2):208–209, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1745576.

Salama:2020:MMI

[SK20] Ibrahim Salama and Gary Koch. On the maximum–minimums identity: Extension and applications. The American Statistician, 74(3):297–300, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1638832.

Siegfried:2023:DFL

[SKH23] Sandra Siegfried, Lucas Kook, and Torsten Hothorn. Distributionfree location-scale regression. The American Statistician, 77(4): 345–356, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2203177.

Sage:2022:BBS

[SLS22] Andrew J. Sage, Yang Liu, and Joe Sato. From black box to shining spotlight: Using random forest prediction intervals to illuminate the impact of assumptions in linear regression. The American Statistician, 76(4):414–429, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2107568.

Sosa:2024:BRS

[Sos24] Juan Sosa. Book review: Statistical Theory: A Concise Introduction, 2nd ed., Felix Abramovich and Ya acov Ritov, Boca Raton, FL: Chapman & Hall/CRC Press, 2023, xiii + 222 pp., $99.95(H), ISBN 978-1-032-00745-8. The American Statistician, 78(3):375, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2024.2325704.

Spitzner:2023:SBR

[Spi23] Dan J. Spitzner. A statistical basis for reporting strength of evidence as pool reduction. The American Statistician, 77(1): 62–71, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2022.2026478.

Singh:2022:REP

[SRSK22] Sushil Kumar Singh, Neelkanth Rawat, Sargun Singh, and Savinder Kaur. Re-exploring the Penney-ante game. The American Statistician, 76(2):177–183, 2022. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2021.
1961860.

Stefan:2022:BRS

[Ste22] Angelika M. Stefan. Book review: Statistics for Making Decisions, by Nicholas T. Longford. Boca Raton, FL: Chapman & Hall/CRC Press, 2021, xv + 292 pp., $120.00 (H), ISBN: 978-0-367-34267-8. The American Statistician, 76(1):87–88, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.2020003.

Sun:2021:LTS

[SW21] Yanlong Sun and Hongbin Wang. Learning temporal structures of random patterns by generating functions. The American Statistician, 75(3):300–309, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1778527.

Swartz:2020:WSP

[Swa20] Tim B. Swartz. Where should I publish my sports paper? The American Statistician, 74(2):103–108, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1459842.

Shi:2021:RVP

[SY21] Haolun Shi and Guosheng Yin. Reconnecting p-value and posterior probability under one- and two-sided tests. The American Statistician, 75(3):265–275, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1717621.

Szabo:2019:TTM

[Sza19] Aniko Szabo. Test for trend with a multinomial outcome. The American Statistician, 73(4):313–320, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2017.

1407823. See comment [Chr20].

Tang:2024:NMC

[Tan24] Yanbo Tang. A note on Monte Carlo integration in high dimensions. The American Statistician, 78(3):290–296, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2267637.

Taraldsen:2020:MMS

[Tar20] Gunnar Taraldsen. Micha Mandel (2020), “The Scaled Uniform Model Revisited,” The American Statistician, 74:1, 98– 100: Comment. The American Statistician, 74(3):315, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1769727. See [Man20].

Tec:2023:CTB

[TDM23] Mauricio Tec, Yunshan Duan, and Peter Müller. A comparative tutorial of Bayesian sequential design and reinforcement learning. The American Statistician, 77(2):223–233, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2022.2129787.

Thomas:2021:LHM

[TT21] Samuel Thomas and Wanzhu Tu. Learning Hamiltonian Monte Carlo in R. The American Statistician, 75(4):403–413, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1865198.

Tang:2021:TPQ

[TWET21] Chuan-Fa Tang, Dewei Wang, Hammou El Barmi, and Joshua M. Tebbs. Testing for positive quadrant dependence. The American Statistician, 75(1):23–30, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2019.
1607554.

VanDerwerken:2021:SPP

[Van21] Douglas VanDerwerken. Slugging percentage is not a percentage — and why that matters. The American Statistician, 75 (2):124–127, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2018.1564698.

vandenHeuvel:2022:MAL

[vdHZ22] Edwin van den Heuvel and Zhuozhao Zhan. Myths about linear and monotonic associations: Pearson’s r, Spearman’s ρ, and Kendall’s τ. The American Statistician, 76(1):44–52, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.2004922.

vanDorp:2020:JSF

[vDJ20] Johan René van Dorp and M. C. Jones. The Johnson system of frequency curves — historical, graphical, and limiting perspectives. The American Statistician, 74(1):37–52, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1637778.

Vexler:2024:CMM

[VH24] Albert Vexler and Alan D. Hutson. A characterization of most (more) powerful test statistics with simple nonparametric applications. The American Statistician, 78(1):36–46, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2023.2192746.

Vedensky:2023:LPP

[VPH23] Daniel Vedensky, Paul A. Parker, and Scott H. Holan. A look into the problem of preferential sampling through the lens of survey statistics. The American Statistician, 77(3):313–322, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2022.2143898.

Vos:2021:LEZ

[VW21] Paul Vos and Qiang Wu. Letter to the Editor: Zhang, J. (2021), “The Mean Relative Entropy: An Invariant Measure of Estimation Error,” The American Statistician, 75, 117–123: comment by Vos and Wu. The American Statistician, 75(4):455– 457, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00031305.2021.1978544. See [Zha21b, Zha21c].

vanZwet:2022:PID

[vZG22] Erik van Zwet and Andrew Gelman. A proposal for informative default priors scaled by the standard error of estimates. The American Statistician, 76(1):1–9, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.1938225.

Weiss:2022:CBP

[WA22] Christian H. Weiß and Boris Aleksandrov. Computing (bivariate) Poisson moments using Stein–Chen identities. The American Statistician, 76(1):10–15, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2020.
1763836.

Waller:2020:GCM

[Wal20] Niels G. Waller. Generating correlation matrices with specified eigenvalues using the method of alternating projections. The American Statistician, 74(1):21–28, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1401960.

Wallin:2024:BRI

[Wal24] Gabriel Wallin. Book review: An Introduction to R and Python for Data Analysis: A Side-by-Side Approach, Taylor R. Brown. Boca Raton, FL: Chapman & Hall/CRC Press, 2023, xix + 246 pp., $99.95(H), ISBN: 978-1-032-20325-6. The American Statistician, 78(2):265, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2024.2320949.

Wang:2023:QDS

[Wan23] Huan Wang. Quantitative drug safety and benefit-risk evaluation: Practical and cross-disciplinary approaches. The American Statistician, 77(1):111–112, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2160592.

Warton:2023:GSE

[War23] David I. Warton. Global simulation envelopes for diagnostic plots in regression models. The American Statistician, 77(4):

425–431, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2022.2139294.

Chang:2024:BAD

[wCW24] Hsin wen Chang and Shu-Hsiang Wang. Bivariate analysis of distribution functions under biased sampling. The American Statistician, 78(2):171–179, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2249965.

Wells:2020:BRS

[Wel20] Jonathan M. Wells. Book review: Surprises in Probability: Seventeen Short Stories, Henk Tijms. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xiii + 131 pp., $29.95(P), ISBN: 978-0-367-00043-1. The American Statistician, 74(3):311, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2020.1790219.

Wells:2021:BRS

[Wel21] Christine R. Wells. Book review: SAS for Mixed Models: Introduction and Basic Applications, by Walter W. Stroup, George A. Milliken, Elizabeth A. Claassen, and Russell D. Wolfinger. Cary, NC: SAS Institute, 2018, xii + 594 pp., $105.95 (paperback), ISBN: 978-1-63526-135-6. The American Statistician, 75(2):231, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2021.1907997.

West:2023:BRA

[Wes23] Brady T. West. Book review: ANOVA and Mixed Models: A Short Introduction Using R. The American Statistician, 77(4): 449–450, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2261817.

Weld:2020:PLR

[WLL20] Christopher Weld, Andrew Loh, and Lawrence Leemis. Plotting likelihood-ratio-based confidence regions for two-parameter univariate probability models. The American Statistician, 74 (2):156–168, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2018.1564696.

Wang:2023:MRP

[WMX^{+}23] Peng Wang, Yilei Ma, Siqi Xu, Yi-Xin Wang, Yu Zhang, Xiangyang Lou, Ming Li, Baolin Wu, Guimin Gao, Ping Yin, and Nianjun Liu. MOVER-R and penalized MOVER-R confidence intervals for the ratio of two quantities. The American Statistician, 77(4):381–389, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2023.2173294.

Wang:2022:CCA

[WSCP22] Wei Wang, Dylan S. Small, Guy Cafri, and Elizabeth W. Paxton. The case-control approach can be more powerful for matched pair observational studies when the outcome is rare. The American Statistician, 76(2):117–123, 2022. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.1972835.

Xiong:2023:SPR

[XC23] Xin Xiong and Ivor Cribben. The state of play of reproducibility in statistics: an empirical analysis. The American Statistician, 77(2):115–126, 2023. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2022.2131625.

Xu:2020:AFR

[XGH^{+}20] Li Xu, Chris Gotwalt, Yili Hong, Caleb B. King, and William Q. Meeker. Applications of the fractional-random-weight bootstrap. The American Statistician, 74(4):345–358, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2020.1731599.

Yang:2021:IAJ

[YK21] Jianning Yang and John E. Kolassa. The impact of application of the jackknife to the sample median. The American Statistician, 75(4):445–449, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2020.1869090.

Young:2021:BRP

[You21] Gabriel J. Young. Book review: Probability and Statistical Inference: From Basic Principles to Advanced Models, by Miltiadis C. Mavrakakis and Jeremy Penzer. Boca Raton, FL: Chapman & Hall/CRC Press, 2021, xix + 423 pp., $99.95 (H/e-book), ISBN: 978-1-58488-939-7 (H), 978-1-315-36663-0 (ebook). The American Statistician, 75(4):451–453, 2021. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.1985863.

Yang:2024:DPM

[YR24] Chengxin Yang and Jerome P. Reiter. Differentially private methods for releasing results of stability analyses. The American Statistician, 78(2):180–191, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2023.
2252870.

Yeh:2022:ERT

[YRD22] Chi-Kuang Yeh, Gregory Rice, and Joel A. Dubin. Evaluating real-time probabilistic forecasts with application to national basketball association outcome prediction. The American Statistician, 76(3):214–223, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1967781.

Yavuz:2020:FUD

[YW20] Fulya Gokalp Yavuz and Mark Daniel Ward. Fostering undergraduate data science. The American Statistician, 74(1):8–16, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1407360.

Zimmermann:2022:PRB

[ZBB^{+}22] Georg Zimmermann, Edgar Brunner, Werner Brannath, Martin Happ, and Arne C. Bathke. Pseudo-ranks: The better way of ranking? The American Statistician, 76(2):124–130, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2021.1972836.

Zimmerman:2022:DSC

[ZH22] Dale L. Zimmerman and Jay M. Ver Hoef. On deconfounding spatial confounding in linear models. The American Statistician, 76(2):159–167, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1946149.

Zhang:2021:FER

[Zha21a] Chunming Zhang. Further examples related to correlations between variables and ranks. The American Statistician, 75(2): 226–229, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1831956.

Zhang:2021:MRE

[Zha21b] Jin Zhang. The mean relative entropy: An invariant measure of estimation error. The American Statistician, 75(2):117–123, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00031305.2018.1543139. See letter [VW21] and re-

sponse [Zha21c].

Zhang:2021:RLE

[Zha21c] Jin Zhang. Response to letter to the Editor: Zhang, J. (2021). The American Statistician, 75(4):458, 2021. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2021.1982557. See [VW21, Zha21b].

Zhang:2024:IPW

[Zha24] Biao Zhang. Inverse probability weighting estimation in completely randomized experiments. The American Statistician, 78 (1):26–35, 2024. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2023.2216247.

Zuo:2022:VSS

[ZSB22] Yi Zuo, Thomas G. Stewart, and Jeffrey D. Blume. Variable selection with second-generation P -values. The American Statistician, 76(2):91–101, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2021.1946150.

Zhang:2020:RFP

[ZZNN20] Haozhe Zhang, Joshua Zimmerman, Dan Nettleton, and Daniel J. Nordman. Random forest prediction intervals. The American Statistician, 74(4):392–406, 2020. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2019.
1585288.

Zimmerman:2021:MMA

[ZZZ21] Dale L. Zimmerman, Nathan D. Zimmerman, and Joshua T. Zimmerman. March madness “Anomalies”: Are they real, and if so, can they be explained? The American Statistician, 75 (2):207–216, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1720814.